### v2

# CELA: Cost-Efficient Language Model Alignment for CTR Prediction

Abstract.
Click-Through Rate (CTR) prediction holds a paramount position in recommender systems. The prevailing ID-based paradigm underperforms in cold-start scenarios due to the skewed distribution of feature frequency. Additionally, the utilization of a single modality fails to exploit the knowledge contained within textual features. Recent efforts have sought to mitigate these challenges by integrating Pre-trained Language Models (PLMs). They design hard prompts to structure raw features into text for each interaction and then apply PLMs for text processing. With external knowledge and reasoning capabilities, PLMs extract valuable information even in cases of sparse interactions. Nevertheless, compared to ID-based models, pure text modeling degrades the efficacy of collaborative filtering, as well as feature scalability and efficiency during both training and inference. To address these issues, we propose Cost-Efficient Language Model Alignment (CELA) for CTR prediction. CELA incorporates textual features and language models while preserving the collaborative filtering capabilities of ID-based models. This model-agnostic framework can be equipped with plug-and-play textual features, with item-level alignment enhancing the utilization of external information while maintaining training and inference efficiency. Through extensive offline experiments, CELA demonstrates superior performance compared to state-of-the-art methods. Furthermore, an online A/B test conducted on an industrial App recommender system showcases its practical effectiveness, solidifying the potential for real-world applications of CELA.
CTR prediction, Pre-trained language model, Textual feature, Cross-Modal Alignment, Recommender Systems
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Recommender systems deliver personalized items to users by analyzing user browsing history and demographics, presenting users with items tailored to their preferences. Click-through Rate (CTR) prediction plays a crucial role in this process by estimating the likelihood that the user will interact with a specific item. This ensures that the displayed items are those most likely to bolster user engagement and revenue.
The ID-based paradigm has garnered considerable success. Notable examples include DeepFM (Guo et al., 2017), PNN (Qu et al., 2016), and DCN (Wang et al., 2017), which focus on high-order explicit feature crossing, while DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) are geared towards mining behavior sequences. This paradigm, illustrated on the left side of Figure 1, entails encoding user and item features into sparse one-hot vectors, subsequently transformed into dense embedding vectors via a lookup process. Latent co-occurrences among features are discerned via a feature interaction layer. Over the past decade, the advancement of these models has centered on developing more sophisticated and scenario-adapted feature interaction layers as a means to enhance the overall performance. Despite their prevalence, ID-based models exhibit certain limitations: 1) Dependence on historical data: This dependence can result in suboptimal modeling in scenarios such as cold-start or long-tail elements, where historical interactions for items or features are limited or absent, leading to inadequate learning of embeddings. 2) Limited feature scope: Primarily focusing on categorical and numerical features, ID-based models overlook other modalities like text, which may harbor untapped knowledge that could potentially enhance the model’s effectiveness.
To alleviate these limitations, works such as PTab (Liu et al., 2022) and CTRL (Li et al., 2023) have integrated Pre-trained Language Models (PLMs) into recommender systems. These models employ hard prompts to structure raw features for each interaction, subsequently processed through PLMs for recommendation purposes, as shown on the right side of the Figure 1. By adopting natural language to represent features instead of one-hot vectors, this approach preserves the semantic integrity of features. This enables PLMs to apply their external knowledge and reasoning capabilities, effectively capturing valuable information even in cases of sparse interactions. For example, ID-based models struggle to learn embeddings for infrequently interacted items like bread, whereas PLMs utilize external knowledge to classify bread as a type of food, akin to toast. Concurrently, this natural language representation facilitates the integration of textual features, such as item descriptions, thereby augmenting the recommender system’s effectiveness.
However, these text-enhanced models face inherent limitations that hinder their practical deployment, also shown in Figure 1. First, modeling based exclusively on pure text may not be ideally suited for capturing collaborative signals, which often results in reduced performance (Li et al., 2023). Second, scalability issues emerge as the number of features increases, potentially leading to constructed text exceeding the language model’s maximum token limit and thus losing valuable information, particularly when the sequence of user behaviors is extensively long. Furthermore, it becomes impractical to delineate user behaviors in a detailed manner, including the specification of item features for each behavior. Last but not least, to bridge the gap between PLM outputs and recommendation tasks, PLMs are usually fine-tuned at the interaction-level (Liu et al., 2022; Geng et al., 2022; Li et al., 2023; Muhamed et al., 2021), yet this process is notably time-intensive. Such inefficiency inherent in PLM text processing extends to inference as well (Liu et al., 2022; Geng et al., 2022; Lin et al., 2023; Bao et al., 2023).
Figure 1. ID-Based Models vs. Pre-trained Language Models (PLMs) for CTR prediction. The former converts tabular data into one-hot vectors, capturing collaborative signals without external knowledge. The latter populates data into a structured prompt, and then feeds the truncated text into a PLM, leveraging external world knowledge for data interpretation.
To overcome these challenges, we propose the Cost-Efficient Language Model Alignment (CELA) for CTR prediction. CELA leverages textual features and PLMs to mitigate the cold-start issue and expand the feature scope. It integrates ID-based models to preserve the collaborative filtering effectiveness, and shifts away from the hard prompt mode, thereby addressing scalability concerns. Through item-level alignment, the PLM is fine-tuned effectively and oriented towards recommendation tasks. CELA is structured as a three-phase paradigm: 1) domain-adaptive pre-training stage, the PLM is further trained on a corpus of textual features, tailoring it to the specific dataset, while also training an ID-based model; 2) recommendation-oriented modal alignment stage, contrastive learning is employed to align the item text representations with the item feature embeddings of the ID-based model. This alignment is at the item level, ensuring minimal training overhead. Post-alignment, text representations of each item are cached, facilitating efficient access via item identifiers, thereby reducing inference costs; and 3) multi-modal feature fusion stage, these aligned text representations are integrated into a new ID-based model, along with non-textual features, enabling their interplay to identify potential click patterns. The final two stages are structured for alternate execution, ensuring ongoing alignment of the PLM with the progressively refined embedding table.
The main contributions of this paper are summarized as follows:
  • We propose a novel paradigm for integrating textual features with ID-based models. This approach is model-agnostic and necessitates minimal modifications to the existing network architecture, rendering it conveniently applicable.
  • We develop an item-level alignment strategy to direct the capabilities of the PLM for recommendation tasks, thereby reducing training time overhead and achieving performance enhancements. Moreover, the cached item text representations are indexed by item identifiers during inference, ensuring low latency.
  • Comprehensive offline experiments on public and industrial datasets demonstrate the effectiveness of CELA. An online A/B test within an App recommender system further validates CELA’s real-world efficacy and feasibility.
2. Related Works
2.1. ID-Based Models for CTR Prediction
ID-based models have been predominant in industry recommendations over the past decade. They capture the characteristics of the items and users by one-hot vectors and learn the collaborative signals to predict the likelihood of a user clicking on an item (Lian et al., 2020). Their architectures are meticulously crafted, evolving into increasingly intricate ones. FM (Rendle, 2010) stands out as one of the earliest machine learning models to incorporate second-order feature interactions. Both WideDeep (Cheng et al., 2016) and DeepFM (Guo et al., 2017) incorporate deep components to learn high-level feature interactions. PNN (Qu et al., 2016) extends the feature interaction function and tightly couples the interaction of first-order and second-order features. AutoInt (Song et al., 2019) and DESTINE (Xu et al., 2021) incorporate self-attention mechanisms to weight different feature interactions. Fi-GNN (Li et al., 2019) leverages graph structures to model feature interactions. DCN (Wang et al., 2017), DCNv2 (Wang et al., 2021), and EDCN (Chen et al., 2021) capture meaningful feature crossings automatically, eliminating the need for manual feature engineering or exhaustive searching. DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) model user behavior sequences, aiming to capture the relationships between the target and historical behaviors.
2.2. Text-Enhanced Models for CTR Prediction
Language models have made remarkable strides, achieving unprecedented natural language understanding and generation. These models are pre-trained on massive datasets and adopt advanced architectures like Transformer (Vaswani et al., 2017), enabling diverse applications such as chatbots (Du et al., 2022), translation (Zhang et al., 2022), and content creation (Chung et al., 2022). Highlighting their potential in semantic comprehension, language models are considered promising for recommendation. Analyzing extra-textual elements such as ad text, search queries, or user reviews, these models may have the capability to discern user intent and subtle nuances between items, thereby enhancing ID-based ones. P5 (Geng et al., 2022) represents a paradigm that redefines various tasks into natural language processing (NLP) tasks. By structuring interactions and metadata in natural language and inputting them into T5 (Raffel et al., 2020), it can generate the target response. CTR-BERT (Muhamed et al., 2021) concatenates all features along with their respective feature names. It undergoes pre-training and distillation with BERT (Devlin et al., 2018). Subsequently, BERT encodes the text of users and items, generating representations that serve as inputs for a late fusion MLP. PTab (Liu et al., 2022) is pre-trained on BERT (Devlin et al., 2018) with a language head using the ”field: value” corpus, which is derived from tabular data. It is then fine-tuned for the CTR task, incorporating a classifier head. UniSRec (Hou et al., 2022) employs a frozen BERT (Devlin et al., 2018) as a text encoder and utilizes a Transformer (Vaswani et al., 2017) architecture to model user behaviors. CTRL (Li et al., 2023) uses contrastive learning for representation alignment between collaborative and semantic models at the interaction level. The collaborative model, now enriched with semantic information, is fine-tuned and can be deployed without the semantic model.
3. Preliminary
3.1. Task Formulation
CTR models are designed to estimate the probability that a user will engage with a specific item—such as music, news, or advertisement—by clicking on it. Mathematically, a CTR model $f$ fits the likelihood $P(\cdot)$ of a click event $y=1$ based on a variety of input features $\boldsymbol{x}$, such as user profile, browsing history, item characteristics, and contextual information. The predicted CTR is formulated as:
(1) $\small\hat{y}=f(\boldsymbol{x}).$
Consider a dataset $\mathcal{D}=\{(\boldsymbol{x}_{k},y_{k})|k=1,\ldots,N\}$, with $N$ as the total number of interactions. Each pair $(\boldsymbol{x}_{k},y_{k})$, where $y_{k}\in\{0,1\}$, represents the $k$-th pair of input features and its corresponding ground truth label. Typically, the model $f$ is trained by minimizing the binary cross-entropy loss:
(2) $\small\mathcal{L}_{BCE}=-\frac{1}{N}\sum^{N}_{k=1}\left[y_{k}\log(\hat{y}_{k})+(1-y_{k})\log(1-\hat{y}_{k})\right].$
3.2. Embedding Layer
Raw input features are usually categorical and then converted into one-hot or multi-hot vectors. In one-hot encoding, each category is represented by a unique binary vector, whose length matches its cardinality $C$, and only one bit is set to 1 to indicate the the presence of that category, as shown in Figure 1. In contrast, multi-hot encoding allows for presence of multiple categories at once.
To compress those sparse, high-dimensional vectors into dense, lower-dimensional ones, the embedding layer is utilized. It acts as a lookup table, denoted by $\mathbf{E}^{(\cdot)}\in\mathbb{R}^{C^{(\cdot)}\times d}$ , which maps each categorical feature into a dense vector of a fixed size $d$. Here, $d$ represents the embedding size. Let $\boldsymbol{x}^{u}$ denote user profile, $\boldsymbol{x}^{h}$ denote browsing history, $\boldsymbol{x}^{c}$ denote contextual information, $\boldsymbol{x}^{i,n}$ denote non-textual features of target item $i$ and $\boldsymbol{x}^{h,n}$ denote non-textual features in user history. This process can be formalized as:
(3) $\boldsymbol{e}=\boldsymbol{x}^{u}\mathbf{E}^{u}\circ\boldsymbol{x}^{h,n}\mathbf{E}^{h}\circ\boldsymbol{x}^{i,n}\mathbf{E}^{i}\circ\boldsymbol{x}^{c}\mathbf{E}^{c},$
where $\circ$ denotes the concatenation function.
4. Methodology
The proposed framework is depicted in Figure 2, encompassing three stages:
  • Domain-Adapted Pre-training (DAP). It involves further pre-training of a PLM on a corpus comprising item-side text specific to the recommendation dataset, to better adapt the PLM to the unique characteristics of the dataset. Concurrently, a traditional ID-based model is trained.
  • Recommendation-Oriented Modal Alignment (ROMA). It focuses on aligning the text representations of items from the PLM with the item-side feature embeddings from the ID-Based model in the latent space. Such alignment is essential for tailoring the text representations to recommendation tasks.
  • Multi-Modal Feature Fusion (MF2). The aligned item text representations are treated as a common feature, which is then integrated with non-textual feature embeddings. Subsequently, a new ID-based model is trained to leverage both the aligned text representations and other features, thereby enhancing the overall effectiveness of the recommender system.
Notably, the last two stages are designed to be executed alternatively, where the PLM is continually aligned to the more advanced embedding table.
Figure 2. The overall framework of CELA. The first stage involves pre-training a PLM on domain-specific item text alongside training an ID-based model. In the second stage, item text representations from the PLM are aligned with the item-side feature embeddings of the ID-based model in latent space. The third stage merges aligned text representations with non-textual features for training a new ID-based model. The final two stages are executed alternately, as denoted by the green dotted line.
4.1. DAP: Domain-Adaptive Pre-training
4.1.1. Language Model Pre-training
To enhance the PLM’s understanding of dataset-specific texts, we construct a corpus from item-side texts, rich in detailed explanations and descriptions that go beyond mere hard prompts and categorical features. Leveraging this corpus for further pre-training, we utilize the Masked Language Model Loss (MLM) (Devlin et al., 2018), which masks a portion of tokens within a sentence and calculates the loss for each masked token prediction. For instance, the input text ”This ballet skirt is nice for dress up play for ballerinas…,” may be masked as ”This ballet skirt is [MASK] for dress-up play for [MASK]…”, forcing the PLM to predict the mask tokens and give the MLM loss as:
(4) $-\𝑚\small\mathcal{L}_{MLM}=-\sum_{i\in\mathcal{I}}\sum_{m}\log P\left(\boldsymbol{x}^{i,t}_{m}|\boldsymbol{x}^{i,t}_{\verb|\|m}\right),$
where $\mathcal{I}$ denotes the item space, $\boldsymbol{x}^{i,t}$ denotes item $i$’s textual feature, $m$ and \$\backslash m$ denote the masked tokens and the rest of the tokens, respectively.
Simultaneously, to enhance item text representations and mitigate the anisotropy issue (Ethayarajh, 2019; Li et al., 2020)—where learned representations are constrained to a narrow cone in vector space, reducing expressiveness—we incorporate SimCSE (Gao et al., 2021) into our framework. SimCSE, an unsupervised technique, employs contrastive learning, treating a text as its own positive pair by feeding it into the PLM twice with standard dropout as the only noise. By drawing positive pairs closer, the PLM is compelled to discern the meanings of texts amidst noise, thereby improving its comprehensive capacity. Moreover, distancing the representations of distinct texts promotes a more uniform distribution of representations in vector space, addressing the anisotropy and yielding superior text representations. This process can be formulated as:
(5) $\small\mathcal{L}_{SimCSE}=-\sum_{i\in\mathcal{I}}\log\frac{\exp{\left(sim_{1}\left(\boldsymbol{h}_{i},\boldsymbol{h}^{\prime}_{i}\right)/\tau_{1}\right)}}{\sum_{j\in\mathcal{I}}\exp\left(sim_{1}\left(\boldsymbol{h}_{i},\boldsymbol{h}^{\prime}_{j}\right)/\tau_{1}\right)},$
where $sim_{1}(\cdot)$ denotes the cosine similarity, $\tau_{1}$ denotes the temperature that scales the similarity measure, and $\boldsymbol{h}_{i}$ and $\boldsymbol{h}^{\prime}_{i}$ denote the representations of item $i$’s text.
The composite loss function for the PLM’s pre-training is:
(6) $\mathcal{L}_{DAP}=\mathcal{L}_{MLM}+\alpha\mathcal{L}_{SimCSE},$
where $\alpha$ is a weighting parameter that balances the contributions of the MLM loss and SimCSE loss. This integrated approach ensures a comprehensive and nuanced adaptation of the PLM to the specifics of the dataset.
4.1.2. ID-Based Model Pre-training
To provide a reference for steering the capabilities of language models towards recommendation tasks, we develop an ID-based model that processes exclusively categorical or numerical features. The model’s prediction function is formulated as:
(7) $\hat{y}=f\left(\boldsymbol{x}^{u},\boldsymbol{x}^{h,n},\boldsymbol{x}^{i,n},\boldsymbol{x}^{c}\right).$
The corresponding training loss is defined in Equation 2.
4.2. ROMA: Recommendation-Oriented Modal Alignment
Despite the profound domain understanding that the PLM achieves through pretraining, a significant gap persists between their output space and that of collaborative filtering. This discrepancy arises because the capabilities of the PLM extend well beyond what recommendation tasks require. Without targeted training and constraints, the outputs of the PLM may not align with the needs of recommender systems. Consequently, it is crucial to refine and orient the PLM’s outputs towards recommendation tasks, a process referred to as alignment. This essential alignment, which aims to provide a representation of the text feature that can be seamlessly integrated with other features for the next stage, is illustrated in the central portion of Figure 2. The substantial number of parameters in the PLM and the extensive volume of interactions in the dataset render interaction-level alignment inefficient (Li et al., 2023; Liu et al., 2022; Geng et al., 2022). Consequently, we propose item-level alignment as a solution.
First, the textual and non-textual features on the item side are encoded into their respective representations. For item $i$, its non-textual features $\boldsymbol{x}^{i,n}$, are encoded utilizing the embedding table from the ID-based model developed in the first stage. Conversely, the textual feature $\boldsymbol{x}^{i,t}$ is processed through the PLM, followed by a transformation via a projection head. This head is to adjust the dimensionality of the output from the PLM to match that of the non-textual features. This encoding process can be formulated as:
(8) $\boldsymbol{e}^{n}_{i}=\boldsymbol{x}^{i,n}\mathbf{E}^{i},$
(9) $\boldsymbol{e}^{t}_{i}=PLM(\boldsymbol{x}^{i,t})\mathbf{W}+\boldsymbol{b},$
where $\boldsymbol{e}^{n}_{i}$ signifies the encoded representation of item $i$’s non-textual features, and $\boldsymbol{e}^{t}_{i}$ represents the encoded textual feature. The terms $\mathbf{W}$ and $\boldsymbol{b}$ are the parameters of the projection head, comprising the matrix and bias, respectively.
Contrastive learning is then utilized to align the representations of textual and non-textual features. Within a given batch, representations of these features from the same item are considered positive pairs, while unmatched representations are treated as negative pairs. Given that both modalities depict the same item, there is an inherent overlap of insights between its positive pair. By incentivizing the model to differentiate between matching and mismatched pairs, contrastive learning facilitates the establishment of correspondence or mapping functions across modalities. The alignment begins with
(10) $\small\mathcal{L}_{t,n}=-\frac{1}{\left|\mathcal{I}\right|}\sum_{i\in\mathcal{I}}\log\frac{\exp\left(sim_{2}\left(\boldsymbol{e}^{t}_{i},\boldsymbol{e}^{n}_{i}\right)/\tau_{2}\right)}{\sum_{k\in\mathcal{I}}\exp\left(sim_{2}\left(\boldsymbol{e}^{t}_{i},\boldsymbol{e}^{n}_{k}\right)/\tau_{2}\right)},$
where $sim_{2}(\cdot)$ denotes the inner product similarity and $\tau_{2}$ denotes another temperature.
To maintain symmetry in the alignment process, a corresponding loss for aligning non-textual representations to textual ones is defined as:
(11) $\small\mathcal{L}_{n,t}=-\frac{1}{\left|\mathcal{I}\right|}\sum_{i\in\mathcal{I}}\log\frac{\exp\left(sim_{2}\left(\boldsymbol{e}^{n}_{i},\boldsymbol{e}^{t}_{i}\right)/\tau_{2}\right)}{\sum_{k\in\mathcal{I}}\exp\left(sim_{2}\left(\boldsymbol{e}^{n}_{i},\boldsymbol{e}^{t}_{k}\right)/\tau_{2}\right)}.$
The overall modal alignment loss combines these two components as:
(12) $\mathcal{L}_{ROMA}=\mathcal{L}_{t,n}+\mathcal{L}_{n,t}.$
Training is exclusively confined to the PLM and projection head, with the embedding table remaining non-trainable to prevent disruption of the collaborative filtering space. It is important to emphasize that, the item-level alignment significantly reduces the computational resources and time required for training.
4.3. MF2: Multi-Modal Feature Fusion
To mitigate suboptimal performance caused by purely text-based recommendations, we preserve the collaborative signals captured by the ID-based model while integrating the PLM as a text encoder. This PLM is tailored to the domain of the textual feature and capable of aligning its outputs to the collaborative space. Serving solely as a text encoder, the PLM prevents scalability issues from increased features, and allows for a detailed description of user behaviors.
To reduce training time and online latency, we create an item text embedding table. Text for each item is processed through the aligned PLM to obtain the embedding, stored for quick access by item identifiers, eliminating the need for real-time PLM processing.
Then, as depicted on the right side of Figure 2, the recommender system is enhanced by integrating the textual feature alongside non-textual features. This textual feature, indexed from the text embedding table using the item’s identifier, is processed using a MLP to ensure dimensional consistency. Within the feature interactions layer, textual and non-textual features are treated uniformly. Such uniformity across all modalities prevents substantial modifications to the original network, thus facilitating its application to any network architecture. This process can be formulated as:
(13) $\boldsymbol{e}^{\prime}=\boldsymbol{e}\circ MLP\left(\boldsymbol{x}^{i,id}~{}\mathbf{E}^{i,t}\right)\circ MLP\left(\boldsymbol{x}^{h,id}~{}\mathbf{E}^{i,t}\right),$
(14) $\hat{y}=FC\left(FI\left(\boldsymbol{e}^{\prime}\right)\right),$
where $\mathbf{E}^{i,t}$ denotes the text embedding table, $\boldsymbol{x}^{i,id}$ denotes the target item’s identifier, $\boldsymbol{x}^{h,id}$ denotes all the items’ identifiers in user behaviors, $FI(\cdot)$ denotes the feature interaction layer and $FC(\cdot)$ denotes the fully connected layer.
4.4. Alternate Training
The last two stages, as indicated by the green dotted line in Figure 2, are designed to be executed in an alternate fashion. Following the completion of the embedding table training in the MF2 stage, we can revert to the ROMA stage. Here, the item text representation is realigned to the newly trained embedding table. Subsequently, this realigned language model is utilized to generate a new text embedding table. This updated table is then employed in the subsequent iteration of the MF2 stage for further training. This alternate approach allows for iterative refinement and enhancement of both the alignment and fusion processes, thereby progressively improving the overall performance of the model.
5. Experiments
In this section, we conduct extensive offline and online experiments to assess CELA by answering these research questions (RQs).
  • RQ1: How does CELA perform against the state-of-the-art?
  • RQ2: How is CELA’s efficiency during training and inference?
  • RQ3: How does CELA perform in cold-start scenarios?
  • RQ4: How do different components influence CELA?
5.1. Experimental Settings
5.1.1. Datasets
We conduct offline experiments on three datasets that include textual features. The statistics for these datasets are provided in Table 1.
[…]
MovieLens Dataset  [^2] https://grouplens.org/datasets/movielens/latest/ collectes the rating activity from MovieLens  [^3] http://movielens.org/. Due to the necessity to extract movie summaries as the textual feature from provided hyperlinks, we sample 10 million interactions through the random sampling of users. The task is to predict whether a user will like the next item, with a preference indicated by a rating of 3 or higher, based on a sequence of up to 20 user behaviors.
AppGallery Dataset comprises application exposure and download records sampled from Huawei AppGallery over eight days. Data from the first seven days constituted the training set, while data from the subsequent day served as the test set. The task is to predict whether the user will download the next application, with application description as the textual feature and a sequence of up to 14 user behaviors.
Table 1. Statistics of datasets for offline experiments.
Dataset #Users #Items #Features #Samples
Amazon 300,562 100,682 6 4,494,230
MovieLens 92,911 26,084 4 10,146,329
AppGallery 11,056,819 29,759 8 16,398,193
5.1.2. Baselines
Our method is benchmarked against contemporary state-of-the-art models, encompassing both ID-based and text-enhanced models. Classic ID-based models, such as AutoInt (Song et al., 2019), DCN (Wang et al., 2017), DCNv2 (Wang et al., 2021), FiBiNET (Huang et al., 2019), PNN (Qu et al., 2016) , WideDeep (Cheng et al., 2016), DeepFM (Guo et al., 2017), and xDeepFM (Lian et al., 2018), focus on modeling high-order feature interactions. In contrast, models like DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) specialize in fine-grained user behavior modeling. Text-enhanced models, including P5 (Geng et al., 2022) and PTab (Liu et al., 2022), aim to describe data in natural language to elicit responses from language models. CTR-BERT (Muhamed et al., 2021) integrates traditional and textual features for fusion in a MLP, UniSRec (Hou et al., 2022) utilizes a frozen BERT (Devlin et al., 2018) as its text encoder, and CTRL (Li et al., 2023) aligns text and tabular data representations to infuse the ID-based model with semantic information.
5.1.3. Evaluation Metrics
In this research, Area Under the Curve (AUC) and logarithmic loss (Logloss) are adopted as evaluation metrics, both widely acknowledged in the assessment of CTR models (Song et al., 2019; Wang et al., 2021; Qu et al., 2016; Guo et al., 2017). AUC evaluates the model’s capacity to accurately rank a positive instance above a negative one, thereby indicating its ranking efficacy. Logloss, on the other hand, quantifies the degree of correspondence between the predicted and actual CTR, offering a metric for the model’s predictive accuracy.
5.1.4. Implementation Details
In our offline experiments, models are implemented using RecStudio (Lian et al., 2023) and Transformers (Wolf et al., 2020). We integrate RoBERTa (Liu et al., 2019) to enhance the optimal ID-based model. AdamW (Loshchilov and Hutter, 2017) optimizer facilitates all training stages. For the DAP stage, $\tau_{1}$ ranges from [0.001 to 0.1] and $\alpha$ from [0.1 to 1.5], with a batch size of 32. In the ROMA stage, $\tau_{2}$ is 1, batch size is 128, the language model’s learning rate is 5e-5, and projection units are [128, $f_{i}d$], with $f_{i}$ as item-side feature count and $d$ set at 16. The MF2 stage uses a batch size of 1024 and standardized hidden units [128, 64, 64] across ID-based models. ID-based models that do not focus on user behavior modeling aggregate user behavior sequences using mean pooling.
5.2. Overall Performance (RQ1)
We conduct a comparative performance evaluation of CELA against baselines, with the results detailed in Table 2. We observe that:
  • Within ID-based models, DIN and DIEN excel on Amazon, because their fine-grained modeling of user behaviors caters to the pronounced reliance on user behaviors. Conversely, for Movielens and AppGallery, the dependency on user behaviors appears less critical, resulting in no marked advantage for DIN and DIEN.
  • CTRL ranks second to CELA in most cases due to its integration of the semantics into the ID-based model, distinguishing it from traditional ID-based models. Unlike other text-enhanced models, CTRL retains collaborative signals, enhancing its performance.
  • Among text-enhanced models, P5 and PTab perform worse than most ID-based models. This observation is consistent with results from (Li et al., 2023), corroborating that recommender systems based purely on text generally underperform compared to ID-based models.
  • CELA demonstrates superior performance across all datasets, thereby validating three core assertions: firstly, the utilization of the PLM extracts intrinsic knowledge within the textual feature; secondly, employing the PLM solely as a textual feature encoder prevents information loss caused by an increase in features and enables the detailed delineation of items; thirdly, the integration of collaborative signals significantly enhances model efficacy. This last point is further supported by the superior outcomes observed in the CTRL and UniSRec compared to other baselines.
Table 2. Performance comparison with baselines. The best and the second best results are in bold and underlined, respectively. denotes the statistical significance of the performance difference with respect to the second-best result at the 0.05 level.
Model Amazon MovieLens AppGallery
AUC Logloss AUC Logloss AUC Logloss
AutoInt 0.8793 0.4384 0.8348 0.3694 0.8426 0.4777
DCN 0.8784 0.4407 0.8352 0.3673 0.8420 0.4787
DCNv2 0.8809 0.4335 0.8377 0.3659 0.8432 0.4771
FiBiNET 0.8792 0.4365 0.8353 0.3647 0.8428 0.4779
PNN 0.8839 0.4314 0.8361 0.3670 0.8431 0.4774
WideDeep 0.8798 0.4383 0.8348 0.3675 0.8422 0.4787
DeepFM 0.8798 0.4378 0.8375 0.3642 0.8434 0.4771
xDeepFM 0.8811 0.4332 0.8367 0.3679 0.8435 0.4775
DIN 0.8949 0.4110 0.8367 0.3626 0.8437 0.4771
DIEN 0.8937 0.4183 0.8343 0.3659 0.8435 0.4816
P5 0.8847 0.4406 0.8197 0.3820 0.8328 0.4920
PTab 0.8811 0.4582 0.8227 0.3809 0.8364 0.4897
CTR-BERT 0.8763 0.4406 0.8294 0.3676 0.8384 0.4846
UniSRec 0.8826 0.4316 0.8381 0.3620 0.8459 0.4736
CTRL 0.8954 0.4100 0.8385 0.3655 0.8464 0.4729
CELA 0.8996* 0.3996* 0.8426* 0.3574* 0.8481* 0.4704*
5.3. Efficiency Measurement (RQ2)
To explore the practicality of CELA in industrial scenarios, we measure its training duration for one epoch and inference latency for a single batch on the industrial dataset AppGallery, compared to baselines. The training duration is calculated for the most time-consuming stage. Results presented in Table 3 reveals that:
  • Within text-enhanced models, the training expenses of P5, PTab, CTR-BERT, and CTRL are notably high, attributable to their complexity, which is implied by the number of parameters. the inference overhead of P5, PTab, and CTR-BERT is also very high for the same reason. Although reducing the update frequency can reduce training costs, employing an outdated language model adversely affects recommendation performance. Therefore, it is not practical to apply them in real-world scenarios.
  • UniSRec, by not fine-tuning the language model, maintains training and inference costs comparable to those of DIN. However, using a frozen language model limits its efficacy in recommendation tasks. This limitation stems from the insufficient alignment of the PLM’s knowledge with the specific requirements of the recommendation domain, resulting in suboptimal integration of textual and non-textual features.
  • The ROMA stage of CELA necessitates fine-tuning of the language model, but this tuning is conducted at the item level rather than the interaction level. Typically, the number of items is significantly lower than that of interactions, leading to minimal training costs, particularly in contexts with a limited number of items. The training and inference costs of CELA’s MF2 are on par with those of DIN, making CELA industry-friendly.
Table 3. Comparison of training and inference efficiency between different models on AppGallery. Params is the number of trainable parameters, the training time is calculated for the whole process and the inference time is calculated for one batch.
Model Params (M) Inference (ms) Total (h)
DIN 0.58 61.38 1.76
P5 60.75 1125.13 170.68
PTab 102.38 256.54 405.72
CTR-BERT 102.98 241.55 96.41
UniSRec 0.54 68.49 3.04
CTRL 102.98 61.38 52.29
CELA DAP (102.88+0.58) 103.46 / (1.07+1.76) 2.83
CELA ROMA 102.30 / 0.30
CELA MF2 0.63 69.64 2.29
CELA / / 5.42
5.4. Performance w.r.t. Long-tail Items (RQ3)
Figure 3. Performance comparison w.r.t. long-tail items on Amazon. The bar graph represents the number of interactions within test data for each group, while the line chart represents the AUC.
To assess CELA’s efficacy in cold-start scenarios, the test dataset is segmented into groups based on item popularity, with the AUC evaluated for each group. As illustrated in Figure 3, CELA consistently outperforms the baseline across all cold-start conditions. This superior performance is attributed to the incorporation of a PLM that leverages embedded external knowledge for text comprehension. Notably, when the popularity falls below 30, the improvements of CELA become particularly pronounced. This is because, under such conditions, ID-based models struggle to learn effective embeddings owing to their dependence on historical interactions.
5.5. Ablation Study (RQ4)
To explore the impact of different training processes and components on the overall performance of CELA, we conduct an ablation study with designed several variants:
  • DIN + RoBERTa: Incorporate RoBERTa as the text encoder and directly undergo the MF2 stage with DIN. It is equivalent to our CELA without the DAP and ROMA stages.
  • CELA w/o DAP: The DAP stage is removed from CELA.
  • CELA w/o ROMA: The ROMA stage is removed from CELA.
  • DAP w/o SimCSE: The SimCSE loss is removed in DAP.
  • DAP w/o MLM: The MLM loss is removed in DAP.
  • ROMA w/o PLM: A text embedding table encoded by the PLM replaces the PLM in ROMA.
  • ROMA w/ cos: Utilizing cosine similarity in ROMA.
  • MF2 w/o text: The text embedding table in MF2 is replaced by DAP’s embedding table from the ID-based model.
Analysis of Table 4 reveals the following insights:
  • DIN + RoBERTa yields fewer performance improvements compared to CELA, suggesting that the mere inclusion of textual feature is insufficient for effective information extraction and utilization in recommendation tasks. The extent of the performance impact varies with the degree of overlap between the PLM’s corpus and the textual features in the dataset.
  • Removing the DAP stage from CELA leads to a reduction in performance, as it deprives the PLM of the opportunity to comprehend domain-specific knowledge, and the anisotropy issue compromises the PLM’s output expressiveness. Nevertheless, the subsequent ROMA stage aligns the PLM output with the recommendation domain, enabling the recommender system to understand PLM outputs. Although the PLM’s knowledge is not perfectly tailored for the domain, it results in performance that still surpasses that of DIN.
  • The removal of the ROMA stage from CELA also leads to diminished performance. While the DAP stage augments the PLM’s capability to comprehend and express the textual feature, a discrepancy between the PLM output space and the collaborative space hinders the recommender system’s understanding of PLM outputs for CTR prediction.
  • Removing the SimCSE loss from DAP leads to a performance decline, as it overlooks the anisotropy issue present in PLMs, thereby diminishing expressiveness. Incorporating SimCSE loss is advisable to enhance the quality of text representations.
  • Removing the MLM loss from DAP results in diminished performance. This is because MLM, serving as a widely adopted pretraining loss, facilitates the swift adaptation of language models to specialized domains. DAP w/o MLM performs even worse than CELA w/o DAP on Amazon dataset, suggesting that using SimCSE loss only in pretraining is not ideal due to the inconsistency with the objective of PLMs’ training. Originally, SimCSE loss was proposed as an auxiliary objective in pretraining, aiming to regularize pre-trained embeddings’ anisotropic space to be more uniform.
  • Removing the PLM significantly reduces performance due to less semantic information preservation when using ROMA directly on the text embedding table. Directly mapping the text embedding table to the ID-based model’s output space, with only a basic projection head, leads to substantial semantic information loss.
  • Replacing the dot product with cosine in ROMA results in reduced performance in AppGallery dataset. So we leverage dot product as the similarity function, and it’s widely used in recommendation system, such as SASRec (Kang and McAuley, 2018), UniSRec (Hou et al., 2022).
  • The performance of MF2 w/o text is comparable to that of DIN, but significantly lags behind CELA. This indicates that CELA’s performance improvement benefits from semantic information.
These observations underscore the importance of both the DAP and ROMA stages in CELA for maximizing the PLM’s potential in recommendation systems, highlighting the critical role of adapting the PLM to domain-specific knowledge and tailoring it for task-specific objectives.
Table 4. Ablation study. The language model is RoBERTa and backbone is DIN.
Variants Amazon AppGallery
AUC Logloss AUC Logloss
DIN 0.8949 0.4110 0.8437 0.4771
DIN + RoBERTa 0.8952 0.4127 0.8472 0.4719
CELA w/o DAP 0.8986 0.4012 0.8475 0.4715
CELA w/o ROMA 0.8965 0.4050 0.8475 0.4709
DAP w/o SimCSE 0.8989 0.4033 0.8478 0.4710
DAP w/o MLM 0.8981 0.4047 0.8477 0.4711
ROMA w/o PLM 0.8669 0.5534 0.8463 0.4731
ROMA w/ cos 0.8997 0.4001 0.8470 0.4715
MF2 w/o text 0.8959 0.4089 0.8440 0.4765
CELA 0.8996 0.3996 0.8481 0.4705
5.6. Further Analysis
5.6.1. Compatibility for language models.
We compare the performance of CELA equipped with PLMs of various sizes: BERTTiny (14.5M) (Jiao et al., 2019), BERTSmall (29.1M) (Bhargava et al., 2021), BERTMedium (41.7M) (Turc et al., 2019), RoBERTaBase (110M) (Liu et al., 2019), BERTLarge (336M) (Devlin et al., 2018), and OPT (1.3B) (Zhang et al., 2022). For the OPT model, Causal Language Modeling (CLM) pre-training loss and LoRA fine-tuning (Hu et al., 2022) are implemented. Experimental results are illustrated in Figure 4, leading to the following insights:
  • CELA, when enhanced with a PLM, no matter its size, exhibits significant performance gains over DIN, validating that the use of PLMs can extract knowledge from textual features and benefit CTR prediction.
  • As the complexity of the comprehensive PLM escalates, the marginal gains in CELA’s performance exhibit a plateauing trend. Notably, BERTLarge underperforms relative to RoBERTaBase, attributable to the former’s significantly larger parameter set, which does not fine-tune as effectively with limited items. This observation implies that beyond a specific threshold, the addition of parameters ceases to yield proportional improvements, suggesting that a medium-sized PLM is sufficient to attain optimal performance.
  • CELA with OPT exhibits a decrease in performance compared to comprehensive PLMs. This reduction is not only because the extensive parameter space of such models proves challenging to fine-tune effectively with a limited item dataset, but it is also likely due to the generative PLM’s unsuitability as a text encoder for CTR prediction.
Figure 4. Performance comparison of CELAs with different PLMs on Amazon. The backbone is indicated by N/A.
5.6.2. Compatibility for ID-based models.
Given the orthogonal nature of our proposed framework to ID-based models, we explore its efficacy across a variety of backbones, such as AutoInt, DCNv2, DeepFM, and DIN. As shown in Table 5, CELA consistently achieves significant performance improvements in comparison to these backbones, thereby confirming the framework’s model-agnostic nature and extensive compatibility.
Table 5. Performance comparison of models with different backbones. The language model is RoBERTa.
Model Amazon AppGallery
AUC Logloss AUC Logloss
AutoInt 0.8793 0.4383 0.8426 0.4777
$\text{CELA}_{\text{AutoInt}}$ 0.8831 0.4323 0.8470 0.4715
DCNv2 0.8809 0.4335 0.8432 0.4771
$\text{CELA}_{\text{DCNv2}}$ 0.8851 0.4290 0.8472 0.4714
DeepFM 0.8798 0.4378 0.8434 0.4771
$\text{CELA}_{\text{DeepFM}}$ 0.8852 0.4263 0.8483 0.4699
DIN 0.8949 0.4110 0.8437 0.4771
$\text{CELA}_{\text{DIN}}$ 0.8996 0.3996 0.8481 0.4705
5.6.3. Alignment of Popular Items Only
ID-based models exhibit poor performance in cold-start scenarios, suggesting inadequacies in learning item embeddings for cold-start items. We aim to ascertain whether aligning poorly learned item embeddings during the ROMA stage might lead to a decline in performance. To this end, we design an experiment in which CELA’s ROMA stage aligns only the top 20%, 40%, 60%, 80%, and 100% of items by popularity. Figure 5 indicate that CELA’s performance reaches its apex when aligning 80% of the most popular items. This suggests that limiting the alignment to a subset of popular items facilitates a more accurate mapping of item text representations to the collaborative filtering space, while aligning poorly learned embeddings will lead to performance degradation.
Figure 5. Performance with alignment of popular items only, evaluated on Amazon. The $x$-axis represents the top percentile of popular items.
5.7. Online A/B Testing
To evaluate the effect of CELA on real-world scenarios, we deploy CELA and conduct an online A/B test in an industrial Android AppGallery, which is one of most popular App stores of China.
5.7.1. Deployment
The online deployment process is outlined as:
  • Data Preprocessing. Utilizing a LLM, we refine the collected app descriptions to eliminate extraneous details, such as contact information, thereby enhancing conciseness. The refined text is indexed by the item identifier for low-cost storage.
  • Update Process. Given the infrequent changes in app descriptions, weekly training is conducted during the DAP stage. Thanks to the efficiency of the ROMA stage, it is feasible to align the PLM daily with the item embeddings from the previous day’s recommendation model. This alignment supplies the MF2 stage with the text embedding table, minimizing additional training overhead. Accordingly, updates to the MF2 stage also occur daily.
  • Online Services. The online recommendation model retrieves text representations directly from the pre-computed text embedding table via item identifiers, circumventing the need for real-time PLM text processing.
This deployment facilitates iterative and daily enhancements of both the ROMA and MF2 stages, while the efficient and effective online serving ensures optimal service delivery.
5.7.2. Online Experimental Results
Table 6. Performance of CELA in online A/B test.
DTR Impr eCPM Impr Latency
0.92% 0.97% 69.73ms
An online A/B test of the CELA is conducted, utilizing 1% of AppGallery’s extensive traffic—which includes tens of millions of daily active users—over a week. The treatment group implements CELA’s three-stage framework, extending the control group.
Table 6 reveals CELA’s significant impact: a 0.92% increase in Download-Through Rate (DTR) and a 0.97% rise in Effective Cost Per Mille (eCPM), indicating notable improvements in user engagement and revenue. Importantly, the online serving latency of CELA, complying with the scenario’s standard of under 100ms and compared to the control group’s 60.04ms, affirms its suitability for real-world applications.
6. Conclusion
In this study, we introduced a novel cost-efficient framework for CTR prediction that leverages the strengths of PLMs and textual features alongside ID-based models. To tailor the capabilities of PLMs for recommendation, we implemented domain-adaptive pre-training and item-level alignment, concurrently addressing the inefficiencies associated with PLM training at the interaction level. By abolishing hard prompts and employing PLMs solely as text encoders, we intricately describe user behaviors while preventing token overflow due to the increase in features. Our framework is model-agnostic and industry-friendly, with low training overhead and serving latency. Comprehensive experiments conducted on both public and industrial datasets validate CELA’s notable enhancements in performance.

### v3

@@
 To address these issues, we propose Cost-Efficient Language Model Alignment (CELA) for CTR prediction.
-CELA incorporates textual features and language models while preserving the collaborative filtering capabilities of ID-based models.
+CELA incorporates item textual features and language models while preserving the collaborative filtering capabilities of ID-based models.
 This model-agnostic framework can be equipped with plug-and-play textual features, with item-level alignment enhancing the utilization of external information while maintaining training and inference efficiency.
 Through extensive offline experiments, CELA demonstrates superior performance compared to state-of-the-art methods.
-Furthermore, an online A/B test conducted on an industrial App recommender system showcases its practical effectiveness, solidifying the potential for real-world applications of CELA.
+Furthermore, an online A/B test conducted on an industrial advertising recommender system showcases its practical effectiveness, solidifying the potential for real-world applications of CELA.
+Codes are available at https://github.com/pepsi2222/CELA.
 CTR Prediction, Pre-trained Language Model, Textual Feature, Cross-Modal Alignment, Recommender Systems
@@
 This ensures that the displayed items are those most likely to bolster user engagement and revenue.
-The ID-based paradigm has garnered considerable success.
-Notable examples include DeepFM (Guo et al., 2017), PNN (Qu et al., 2016), and DCN (Wang et al., 2017), which focus on high-order explicit feature crossing, while DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) are geared towards mining behavior sequences.
-This paradigm, illustrated on the left side of Figure 1, entails encoding user and item features into sparse one-hot vectors, subsequently transformed into dense embedding vectors via a lookup process.
-Latent co-occurrences among features are discerned via a feature interaction layer.
-Over the past decade, the advancement of these models has centered on developing more sophisticated and scenario-adapted feature interaction layers as a means to enhance the overall performance.
-Despite their prevalence, ID-based models exhibit certain limitations: 1) Dependence on historical data: This dependence can result in suboptimal modeling in scenarios such as cold-start or long-tail elements, where historical interactions for items or features are limited or absent, leading to inadequate learning of embeddings.
-2) Limited feature scope: Primarily focusing on categorical and numerical features, ID-based models overlook other modalities like text, which may harbor untapped knowledge that could potentially enhance the model’s effectiveness.
-To alleviate these limitations, works such as PTab (Liu et al., 2022) and CTRL (Li et al., 2023) have integrated Pre-trained Language Models (PLMs) into recommender systems.
-These models employ hard prompts to structure raw features for each interaction, subsequently processed through PLMs for recommendation purposes, as shown on the right side of the Figure 1.
-By adopting natural language to represent features instead of one-hot vectors, this approach preserves the semantic integrity of features.
-This enables PLMs to apply their external knowledge and reasoning capabilities, effectively capturing valuable information even in cases of sparse interactions.
-For example, ID-based models struggle to learn embeddings for infrequently interacted items like bread, whereas PLMs utilize external knowledge to classify bread as a type of food, akin to toast.
-Concurrently, this natural language representation facilitates the integration of textual features, such as item descriptions, thereby augmenting the recommender system’s effectiveness.
-However, these text-enhanced models face inherent limitations that hinder their practical deployment, also shown in Figure 1.
-First, modeling based exclusively on pure text may not be ideally suited for capturing collaborative signals, which often results in reduced performance (Li et al., 2023).
-Second, scalability issues emerge as the number of features increases, potentially leading to constructed text exceeding the language model’s maximum token limit and thus losing valuable information, particularly when the sequence of user behaviors is extensively long.
-Furthermore, it becomes impractical to delineate user behaviors in a detailed manner, including the specification of item features for each behavior.
-Last but not least, to bridge the gap between PLM outputs and recommendation tasks, PLMs are usually fine-tuned at the interaction-level (Liu et al., 2022; Geng et al., 2022; Li et al., 2023; Muhamed et al., 2021), yet this process is notably time-intensive.
-Such inefficiency inherent in PLM text processing extends to inference as well (Liu et al., 2022; Geng et al., 2022; Lin et al., 2023; Bao et al., 2023).
 Figure 1.
@@
 The former converts tabular data into one-hot vectors, capturing collaborative signals without external knowledge.
-The latter populates data into a structured prompt, and then feeds the truncated text into a PLM, leveraging external world knowledge for data interpretation.
+The latter populates data into a hard prompt, and then feeds truncated texts into the PLM, leveraging external world knowledge for data interpretation.
+The ID-based paradigm has garnered considerable success.
+Notable examples include DeepFM (Guo et al., 2017), PNN (Qu et al., 2016), and DCN (Wang et al., 2017), which focus on high-order explicit feature crossing, while DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) are geared towards mining behavior sequences.
+This paradigm, illustrated on the left side of Figure 1, entails encoding user and item features into sparse one-hot vectors, subsequently transforming into dense embedding vectors via a lookup process.
+Latent co-occurrences among features are discerned via a feature interaction layer.
+Over the past decade, the advancement of these models has centered on developing more sophisticated and scenario-adapted feature interaction layers as a means to enhance overall performance.
+Despite their prevalence, ID-based models exhibit certain limitations: 1) Dependence on historical data: This dependence can result in suboptimal modeling in scenarios such as cold-start or long-tail elements, where historical interactions for items or features are limited or absent, leading to inadequate learning of embeddings.
+2) Limited feature scope: Primarily focusing on categorical and numerical features, ID-based models overlook other modalities like text, which may harbor untapped knowledge that could potentially enhance the model’s effectiveness.
+To alleviate these limitations, works such as PTab (Liu et al., 2022) and CTRL (Li et al., 2023) have integrated Pre-trained Language Models (PLMs) into recommender systems.
+These models employ hard prompts to structure raw features for each interaction, subsequently processed through PLMs for recommendation purposes, as shown on the right side of Figure 1.
+By adopting natural language to represent features instead of one-hot vectors, this approach preserves the semantic integrity of features.
+This enables PLMs to apply their external knowledge and reasoning capabilities, effectively capturing valuable information even in cases of sparse interactions.
+For example, ID-based models struggle to learn embeddings for infrequently interacted items like bread, whereas PLMs utilize external knowledge to classify rice as a type of food, akin to toast.
+Concurrently, this natural language representation facilitates the integration of textual features, such as item descriptions, thereby augmenting the recommender system’s effectiveness.
+However, these text-enhanced models face inherent limitations that hinder their practical deployment, shown in Figure 1.
+First, modeling based exclusively on pure text may not be ideally suited for capturing collaborative signals, which often results in reduced performance (Li et al., 2023).
+Second, scalability issues emerge as the number of features increases, potentially leading to constructed text exceeding the language model’s maximum token limit and thus losing valuable information, particularly when the sequence of user behaviors is extensively long.
+Furthermore, it becomes impractical to delineate user behaviors in a detailed manner, including the specification of item features for each behavior (Yang et al., 2024).
+Last but not least, to bridge the gap between PLM outputs and recommendation tasks, PLMs are usually fine-tuned at the interaction-level (Liu et al., 2022; Geng et al., 2022; Li et al., 2023; Muhamed et al., 2021), yet this process is notably time-intensive.
+Such inefficiency inherent in PLM text processing extends to inference as well (Liu et al., 2022; Geng et al., 2022; Lin et al., 2024; Bao et al., 2023).
 To overcome these challenges, we propose the Cost-Efficient Language Model Alignment (CELA) for CTR prediction.
-CELA leverages textual features and PLMs to mitigate the cold-start issue and expand the feature scope.
-It integrates ID-based models to preserve the collaborative filtering effectiveness, and shifts away from the hard prompt mode, thereby addressing scalability concerns.
-Through item-level alignment, the PLM is fine-tuned effectively and oriented towards recommendation tasks.
-CELA is structured as a three-phase paradigm: 1) domain-adaptive pre-training stage, the PLM is further trained on a corpus of textual features, tailoring it to the specific dataset, while also training an ID-based model; 2) recommendation-oriented modal alignment stage, contrastive learning is employed to align the item text representations with the item feature embeddings of the ID-based model.
-This alignment is at the item level, ensuring minimal training overhead.
-Post-alignment, text representations of each item are cached, facilitating efficient access via item identifiers, thereby reducing inference costs; and 3) multi-modal feature fusion stage, these aligned text representations are integrated into a new ID-based model, along with non-textual features, enabling their interplay to identify potential click patterns.
+To mitigate cold-start issues, CELA expands its feature scope by leveraging item text features and PLMs to generate more accurate item and user profiles.
+To maintain optimality and reduce inference latency, CELA integrates ID-based models, sustaining the efficacy and efficiency of collaborative filtering.
+For scalability, CELA shifts from the hard prompt mode to using PLMs solely as encoders for item textual features.
+Furthermore, to reduce training costs, CELA fine-tunes PLMs through item-level alignment.
+CELA is structured as a three-phase paradigm: 1) Domain-Adaptive Pre-training stage, the PLM is further pre-trained on a corpus of item textual features, tailoring it to the specific dataset; 2) Recommendation-Oriented Modal Alignment stage, contrastive learning is employed to align the PLM with the item ID embedding table of a developed ID-based model.
+This alignment is at the item level, ensuring minimal training overhead; and 3) Multi-Modal Feature Fusion stage, aligned text representations of each item are cached, facilitating efficient access via item identifiers, thereby reducing inference costs.
+Then these text representations are integrated into a new ID-based model, along with non-textual features, enabling their interplay to identify potential click patterns.
 The final two stages are structured for alternate execution, ensuring ongoing alignment of the PLM with the progressively refined embedding table.
 The main contributions of this paper are summarized as follows:
-  • We propose a novel paradigm for integrating textual features with ID-based models.
+  • We propose a novel paradigm for integrating item textual features with ID-based models.
 This approach is model-agnostic and necessitates minimal modifications to the existing network architecture, rendering it conveniently applicable.
@@
   • Comprehensive offline experiments on public and industrial datasets demonstrate the effectiveness of CELA.
-An online A/B test within an App recommender system further validates CELA’s real-world efficacy and feasibility.
+An online A/B test within an industrial advertising recommender system further validates CELA’s real-world efficacy and feasibility.
 2. Related Works
@@
 AutoInt (Song et al., 2019) and DESTINE (Xu et al., 2021) incorporate self-attention mechanisms to weight different feature interactions.
-Fi-GNN (Li et al., 2019) leverages graph structures to model feature interactions.
 DCN (Wang et al., 2017), DCNv2 (Wang et al., 2021), and EDCN (Chen et al., 2021) capture meaningful feature crossings automatically, eliminating the need for manual feature engineering or exhaustive searching.
-DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) model user behavior sequences, aiming to capture the relationships between the target and historical behaviors.
+DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) model user behavior sequences, aiming to capture the relationships between the target item and historical behaviors.
 2.2. Text-Enhanced Models for CTR Prediction
 Language models have made remarkable strides, achieving unprecedented natural language understanding and generation.
-These models are pre-trained on massive datasets and adopt advanced architectures like Transformer (Vaswani et al., 2017), enabling diverse applications such as chatbots (Du et al., 2022), translation (Zhang et al., 2022), and content creation (Chung et al., 2022).
+Pre-trained on extensive datasets and utilizing sophisticated architectures like Transformer (Vaswani et al., 2017), these models support diverse applications such as chatbots (Du et al., 2022), translation (Zhang et al., 2022), and content creation (Chung et al., 2022).
 Highlighting their potential in semantic comprehension, language models are considered promising for recommendation.
-Analyzing extra-textual elements such as ad text, search queries, or user reviews, these models may have the capability to discern user intent and subtle nuances between items, thereby enhancing ID-based ones.
-P5 (Geng et al., 2022) represents a paradigm that redefines various tasks into natural language processing (NLP) tasks.
-By structuring interactions and metadata in natural language and inputting them into T5 (Raffel et al., 2020), it can generate the target response.
-CTR-BERT (Muhamed et al., 2021) concatenates all features along with their respective feature names.
-It undergoes pre-training and distillation with BERT (Devlin et al., 2018).
-Subsequently, BERT encodes the text of users and items, generating representations that serve as inputs for a late fusion MLP.
-PTab (Liu et al., 2022) is pre-trained on BERT (Devlin et al., 2018) with a language head using the ”field: value” corpus, which is derived from tabular data.
-It is then fine-tuned for the CTR task, incorporating a classifier head.
-UniSRec (Hou et al., 2022) employs a frozen BERT (Devlin et al., 2018) as a text encoder and utilizes a Transformer (Vaswani et al., 2017) architecture to model user behaviors.
-CTRL (Li et al., 2023) uses contrastive learning for representation alignment between collaborative and semantic models at the interaction level.
-The collaborative model, now enriched with semantic information, is fine-tuned and can be deployed without the semantic model.
+Analyzing extra-textual elements such as product descriptions, these models may have the capability to discern user intent and item nuances, thereby enhancing ID-based ones.
+P5 (Geng et al., 2022) redefines tasks as natural language processing (NLP) tasks, structuring interactions and metadata in natural language and inputting them into T5 (Raffel et al., 2020) to generate target responses.
+CTR-BERT (Muhamed et al., 2021) concatenates all features with their names, undergoes pre-training and distillation with BERT (Devlin et al., 2019), and encodes user and item text to generate representations for a late fusion MLP.
+PTab (Liu et al., 2022) is pre-trained on BERT using a ”field: value” corpus from tabular data, then fine-tuned for the CTR task with a classifier head.
+UniSRec (Hou et al., 2022) employs a frozen BERT as a text encoder and utilizes a Transformer architecture to model user behaviors.
+CTRL (Li et al., 2023) uses contrastive learning to align representations between collaborative and semantic models at the interaction level.
+The collaborative model, now enriched with semantic information, is fine-tuned and can be deployed independently.
 3. Preliminary
@@
 Here, $d$ represents the embedding size.
-Let $\boldsymbol{x}^{u}$ denote user profile, $\boldsymbol{x}^{h}$ denote browsing history, $\boldsymbol{x}^{c}$ denote contextual information, $\boldsymbol{x}^{i,n}$ denote non-textual features of target item $i$ and $\boldsymbol{x}^{h,n}$ denote non-textual features in user history.
+Let $\boldsymbol{x}^{u}$ denote user profile, $\boldsymbol{x}^{h}$ denote browsing history, $\boldsymbol{x}^{c}$ denote contextual information, $\boldsymbol{x}^{i,n}$ denote non-textual features of candidate item $i$ and $\boldsymbol{x}^{h,n}$ denote non-textual features in user history.
 This process can be formalized as:
@@
 4. Methodology
+Figure 2.
+The overall framework of CELA.
+The first stage pre-trains a PLM on domain-specific item texts.
+In the second stage, an ID-based model is developed and item text representations from the PLM are aligned with the item-side feature embeddings of the ID-based model in latent space.
+The third stage merges aligned text representations with non-textual features for training a new ID-based model.
+The final two stages are executed alternately, as denoted by the green dotted line.
 The proposed framework is depicted in Figure 2, encompassing three stages:
@@
 It involves further pre-training of a PLM on a corpus comprising item-side text specific to the recommendation dataset, to better adapt the PLM to the unique characteristics of the dataset.
-Concurrently, a traditional ID-based model is trained.
   • Recommendation-Oriented Modal Alignment (ROMA).
@@
 Subsequently, a new ID-based model is trained to leverage both the aligned text representations and other features, thereby enhancing the overall effectiveness of the recommender system.
-Notably, the last two stages are designed to be executed alternatively, where the PLM is continually aligned to the more advanced embedding table.
-Figure 2.
-The overall framework of CELA.
-The first stage involves pre-training a PLM on domain-specific item text alongside training an ID-based model.
-In the second stage, item text representations from the PLM are aligned with the item-side feature embeddings of the ID-based model in latent space.
-The third stage merges aligned text representations with non-textual features for training a new ID-based model.
-The final two stages are executed alternately, as denoted by the green dotted line.
+Notably, the last two stages can be executed alternatively, where the PLM is continually aligned to the more advanced ID embedding table.
 4.1. DAP: Domain-Adaptive Pre-training
-4.1.1. Language Model Pre-training
-To enhance the PLM’s understanding of dataset-specific texts, we construct a corpus from item-side texts, rich in detailed explanations and descriptions that go beyond mere hard prompts and categorical features.
-Leveraging this corpus for further pre-training, we utilize the Masked Language Model Loss (MLM) (Devlin et al., 2018), which masks a portion of tokens within a sentence and calculates the loss for each masked token prediction.
-For instance, the input text ”This ballet skirt is nice for dress up play for ballerinas…,” may be masked as ”This ballet skirt is [MASK] for dress-up play for [MASK]…”, forcing the PLM to predict the mask tokens and give the MLM loss as:
+To enhance the PLM’s understanding of dataset-specific texts, we construct a corpus from item-side texts, rich in detailed item descriptions, as illustrated in Appendix A.
+Leveraging this corpus for further pre-training, we utilize the Masked Language Model Loss (MLM) (Devlin et al., 2019), which masks a portion of tokens within a sentence and calculates the loss for each masked token prediction.
+For instance, the input text ”This ballet skirt is nice for dress-up play for ballerinas…,” may be masked as ”This ballet skirt is [MASK] for dress-up play for [MASK]…”, forcing the PLM to predict the mask tokens and give the MLM loss as:
 (4) $-\𝑚\small\mathcal{L}_{MLM}=-\sum_{i\in\mathcal{I}}\sum_{m}\log P\left(\boldsymbol{x}^{i,t}_{m}|\boldsymbol{x}^{i,t}_{\verb|\|m}\right),$
@@
 Simultaneously, to enhance item text representations and mitigate the anisotropy issue (Ethayarajh, 2019; Li et al., 2020)—where learned representations are constrained to a narrow cone in vector space, reducing expressiveness—we incorporate SimCSE (Gao et al., 2021) into our framework.
-SimCSE, an unsupervised technique, employs contrastive learning, treating a text as its own positive pair by feeding it into the PLM twice with standard dropout as the only noise.
+SimCSE, an unsupervised technique, employs contrastive learning, treating a text as its positive pair by feeding it into the PLM twice with standard dropout as the only noise.
 By drawing positive pairs closer, the PLM is compelled to discern the meanings of texts amidst noise, thereby improving its comprehensive capacity.
@@
 This process can be formulated as:
-(5) $\small\mathcal{L}_{SimCSE}=-\sum_{i\in\mathcal{I}}\log\frac{\exp{\left(sim_{1}\left(\boldsymbol{h}_{i},\boldsymbol{h}^{\prime}_{i}\right)/\tau_{1}\right)}}{\sum_{j\in\mathcal{I}}\exp\left(sim_{1}\left(\boldsymbol{h}_{i},\boldsymbol{h}^{\prime}_{j}\right)/\tau_{1}\right)},$
-where $sim_{1}(\cdot)$ denotes the cosine similarity, $\tau_{1}$ denotes the temperature that scales the similarity measure, and $\boldsymbol{h}_{i}$ and $\boldsymbol{h}^{\prime}_{i}$ denote the representations of item $i$’s text.
+(5) $\small\mathcal{L}_{SimCSE}=-\sum_{i\in\mathcal{I}}\log\frac{\exp{\left(sim\left(\boldsymbol{h}_{i},\boldsymbol{h}^{\prime}_{i}\right)/\tau_{1}\right)}}{\sum_{j\in\mathcal{I}}\exp\left(sim\left(\boldsymbol{h}_{i},\boldsymbol{h}^{\prime}_{j}\right)/\tau_{1}\right)},$
+where $sim(\cdot)$ denotes the similarity function, $\tau_{1}$ denotes the temperature scaling the similarity measure, and $\boldsymbol{h}_{i}$ and $\boldsymbol{h}^{\prime}_{i}$ denote the representations of item $i$’s text.
 The composite loss function for the PLM’s pre-training is:
@@
 This integrated approach ensures a comprehensive and nuanced adaptation of the PLM to the specifics of the dataset.
-4.1.2. ID-Based Model Pre-training
-To provide a reference for steering the capabilities of language models towards recommendation tasks, we develop an ID-based model that processes exclusively categorical or numerical features.
-The model’s prediction function is formulated as:
-(7) $\hat{y}=f\left(\boldsymbol{x}^{u},\boldsymbol{x}^{h,n},\boldsymbol{x}^{i,n},\boldsymbol{x}^{c}\right).$
-The corresponding training loss is defined in Equation 2.
 4.2. ROMA: Recommendation-Oriented Modal Alignment
@@
 Consequently, it is crucial to refine and orient the PLM’s outputs towards recommendation tasks, a process referred to as alignment.
-This essential alignment, which aims to provide a representation of the text feature that can be seamlessly integrated with other features for the next stage, is illustrated in the central portion of Figure 2.
+This essential alignment, which aims to provide a representation of the item’s textual feature that can be seamlessly integrated with other features for the next stage, is illustrated in the central portion of Figure 2.
 The substantial number of parameters in the PLM and the extensive volume of interactions in the dataset render interaction-level alignment inefficient (Li et al., 2023; Liu et al., 2022; Geng et al., 2022).
 Consequently, we propose item-level alignment as a solution.
+4.2.1. ID-Based Model
+To provide a reference for steering capabilities of the PLM towards recommendation tasks, we develop an ID-based model that processes exclusively categorical or numerical features.
+The model’s prediction function is formulated as:
+(7) $\hat{y}=f\left(\boldsymbol{x}^{u},\boldsymbol{x}^{h,n},\boldsymbol{x}^{i,n},\boldsymbol{x}^{c}\right).$
+The corresponding training loss is defined in Equation 2.
+4.2.2. Item-Level Modal Alignment
 First, the textual and non-textual features on the item side are encoded into their respective representations.
-For item $i$, its non-textual features $\boldsymbol{x}^{i,n}$, are encoded utilizing the embedding table from the ID-based model developed in the first stage.
+For item $i$, its non-textual features $\boldsymbol{x}^{i,n}$, are encoded utilizing the frozen ID embedding table developed in Section 4.2.1.
 Conversely, the textual feature $\boldsymbol{x}^{i,t}$ is processed through the PLM, followed by a transformation via a projection head.
 This head is to adjust the dimensionality of the output from the PLM to match that of the non-textual features.
-This encoding process can be formulated as:
+This encoding process is formulated as:
 (8) $\boldsymbol{e}^{n}_{i}=\boldsymbol{x}^{i,n}\mathbf{E}^{i},$
 (9) $\boldsymbol{e}^{t}_{i}=PLM(\boldsymbol{x}^{i,t})\mathbf{W}+\boldsymbol{b},$
-where $\boldsymbol{e}^{n}_{i}$ signifies the encoded representation of item $i$’s non-textual features, and $\boldsymbol{e}^{t}_{i}$ represents the encoded textual feature.
-The terms $\mathbf{W}$ and $\boldsymbol{b}$ are the parameters of the projection head, comprising the matrix and bias, respectively.
-Contrastive learning is then utilized to align the representations of textual and non-textual features.
+where $\boldsymbol{e}^{n}_{i}$ represents the encoded representation of item $i$’s non-textual features, and $\boldsymbol{e}^{t}_{i}$ represents the encoded textual feature.
+The terms $\mathbf{W}$ and $\boldsymbol{b}$ are the parameters of the projection head, representing the matrix and bias, respectively.
+Contrastive learning is then utilized to align the representations of textual and non-textual features on the item side.
 Within a given batch, representations of these features from the same item are considered positive pairs, while unmatched representations are treated as negative pairs.
-Given that both modalities depict the same item, there is an inherent overlap of insights between its positive pair.
+Given that both modalities depict the same item, there is an inherent overlap of insights between its positive pair, analogous to the relationship among visual, textual, and auditory data (Girdhar et al., 2023).
 By incentivizing the model to differentiate between matching and mismatched pairs, contrastive learning facilitates the establishment of correspondence or mapping functions across modalities.
 The alignment begins with
-(10) $\small\mathcal{L}_{t,n}=-\frac{1}{\left|\mathcal{I}\right|}\sum_{i\in\mathcal{I}}\log\frac{\exp\left(sim_{2}\left(\boldsymbol{e}^{t}_{i},\boldsymbol{e}^{n}_{i}\right)/\tau_{2}\right)}{\sum_{k\in\mathcal{I}}\exp\left(sim_{2}\left(\boldsymbol{e}^{t}_{i},\boldsymbol{e}^{n}_{k}\right)/\tau_{2}\right)},$
-where $sim_{2}(\cdot)$ denotes the inner product similarity and $\tau_{2}$ denotes another temperature.
+(10) $\small\mathcal{L}_{t,n}=-\frac{1}{\left|\mathcal{I}\right|}\sum_{i\in\mathcal{I}}\log\frac{\exp\left(sim\left(\boldsymbol{e}^{t}_{i},\boldsymbol{e}^{n}_{i}\right)/\tau_{2}\right)}{\sum_{k\in\mathcal{I}}\exp\left(sim\left(\boldsymbol{e}^{t}_{i},\boldsymbol{e}^{n}_{k}\right)/\tau_{2}\right)},$
+where $\tau_{2}$ denotes another temperature scaling the similarity.
 To maintain symmetry in the alignment process, a corresponding loss for aligning non-textual representations to textual ones is defined as:
-(11) $\small\mathcal{L}_{n,t}=-\frac{1}{\left|\mathcal{I}\right|}\sum_{i\in\mathcal{I}}\log\frac{\exp\left(sim_{2}\left(\boldsymbol{e}^{n}_{i},\boldsymbol{e}^{t}_{i}\right)/\tau_{2}\right)}{\sum_{k\in\mathcal{I}}\exp\left(sim_{2}\left(\boldsymbol{e}^{n}_{i},\boldsymbol{e}^{t}_{k}\right)/\tau_{2}\right)}.$
-The overall modal alignment loss combines these two components as:
+(11) $\small\mathcal{L}_{n,t}=-\frac{1}{\left|\mathcal{I}\right|}\sum_{i\in\mathcal{I}}\log\frac{\exp\left(sim\left(\boldsymbol{e}^{n}_{i},\boldsymbol{e}^{t}_{i}\right)/\tau_{2}\right)}{\sum_{k\in\mathcal{I}}\exp\left(sim\left(\boldsymbol{e}^{n}_{i},\boldsymbol{e}^{t}_{k}\right)/\tau_{2}\right)}.$
+The overall modal alignment loss combines these two components:
 (12) $\mathcal{L}_{ROMA}=\mathcal{L}_{t,n}+\mathcal{L}_{n,t}.$
-Training is exclusively confined to the PLM and projection head, with the embedding table remaining non-trainable to prevent disruption of the collaborative filtering space.
+Training is exclusively confined to the PLM and projection head, with the ID embedding table remaining non-trainable to prevent disruption of the collaborative filtering space.
 It is important to emphasize that, the item-level alignment significantly reduces the computational resources and time required for training.
 4.3. MF2: Multi-Modal Feature Fusion
-To mitigate suboptimal performance caused by purely text-based recommendations, we preserve the collaborative signals captured by the ID-based model while integrating the PLM as a text encoder.
-This PLM is tailored to the domain of the textual feature and capable of aligning its outputs to the collaborative space.
-Serving solely as a text encoder, the PLM prevents scalability issues from increased features, and allows for a detailed description of user behaviors.
+To mitigate suboptimal performance caused by purely text-based models, we preserve collaborative signals captured by the ID-based model while integrating the aligned PLM.
+Serving solely as an item textual feature encoder, the PLM avoids scalability issues from increased features and prevents information loss by encoding textual features of user behaviors item by item.
 To reduce training time and online latency, we create an item text embedding table.
-Text for each item is processed through the aligned PLM to obtain the embedding, stored for quick access by item identifiers, eliminating the need for real-time PLM processing.
-Then, as depicted on the right side of Figure 2, the recommender system is enhanced by integrating the textual feature alongside non-textual features.
-This textual feature, indexed from the text embedding table using the item’s identifier, is processed using a MLP to ensure dimensional consistency.
+Text for each item is processed through the aligned PLM to obtain the embedding, and stored for quick access by item identifiers, eliminating the need for real-time PLM processing.
+Then, as depicted on the right side of Figure 2, the recommender system is enhanced by integrating the item’s textual feature alongside non-textual features.
+This item textual feature, indexed from the text embedding table using the item’s identifier, is processed using an MLP to ensure dimensional consistency.
 Within the feature interactions layer, textual and non-textual features are treated uniformly.
 Such uniformity across all modalities prevents substantial modifications to the original network, thus facilitating its application to any network architecture.
-This process can be formulated as:
+This process is formulated as:
 (13) $\boldsymbol{e}^{\prime}=\boldsymbol{e}\circ MLP\left(\boldsymbol{x}^{i,id}~{}\mathbf{E}^{i,t}\right)\circ MLP\left(\boldsymbol{x}^{h,id}~{}\mathbf{E}^{i,t}\right),$
 (14) $\hat{y}=FC\left(FI\left(\boldsymbol{e}^{\prime}\right)\right),$
-where $\mathbf{E}^{i,t}$ denotes the text embedding table, $\boldsymbol{x}^{i,id}$ denotes the target item’s identifier, $\boldsymbol{x}^{h,id}$ denotes all the items’ identifiers in user behaviors, $FI(\cdot)$ denotes the feature interaction layer and $FC(\cdot)$ denotes the fully connected layer.
+where $\mathbf{E}^{i,t}$ denotes the text embedding table, $\boldsymbol{x}^{i,id}$ denotes the candidate item’s ID, and $\boldsymbol{x}^{h,id}$ denotes items’ IDs in user history.
+$\boldsymbol{e}^{\prime}$ is an augmented version of $\boldsymbol{e}$, concatenated with textual embeddings of the candidate item and user history.
+$FI(\cdot)$ denotes the feature interaction layer and $FC(\cdot)$ denotes the fully connected layer.
 4.4. Alternate Training
-The last two stages, as indicated by the green dotted line in Figure 2, are designed to be executed in an alternate fashion.
-Following the completion of the embedding table training in the MF2 stage, we can revert to the ROMA stage.
-Here, the item text representation is realigned to the newly trained embedding table.
-Subsequently, this realigned language model is utilized to generate a new text embedding table.
+The last two stages, as indicated by the green dotted line in Figure 2, are designed to be executed alternately.
+Following the completion of the ID embedding table training in the MF2 stage, we can revert to the ROMA stage.
+Here, the item text representation is realigned to the newly trained ID embedding table.
+Subsequently, this realigned PLM is utilized to generate a new text embedding table.
 This updated table is then employed in the subsequent iteration of the MF2 stage for further training.
-This alternate approach allows for iterative refinement and enhancement of both the alignment and fusion processes, thereby progressively improving the overall performance of the model.
+This alternate approach allows for iterative refinement of both the alignment and fusion processes, thereby progressively improving the overall performance.
 5. Experiments
@@
 5.1.1. Datasets
-We conduct offline experiments on three datasets that include textual features. The statistics for these datasets are provided in Table 1.
-[…]
-MovieLens Dataset  [^2] https://grouplens.org/datasets/movielens/latest/ collectes the rating activity from MovieLens  [^3] http://movielens.org/.
-Due to the necessity to extract movie summaries as the textual feature from provided hyperlinks, we sample 10 million interactions through the random sampling of users.
-The task is to predict whether a user will like the next item, with a preference indicated by a rating of 3 or higher, based on a sequence of up to 20 user behaviors.
-AppGallery Dataset comprises application exposure and download records sampled from Huawei AppGallery over eight days.
-Data from the first seven days constituted the training set, while data from the subsequent day served as the test set.
-The task is to predict whether the user will download the next application, with application description as the textual feature and a sequence of up to 14 user behaviors.
-Table 1. Statistics of datasets for offline experiments.
-Dataset #Users #Items #Features #Samples
-Amazon 300,562 100,682 6 4,494,230
-MovieLens 92,911 26,084 4 10,146,329
-AppGallery 11,056,819 29,759 8 16,398,193
+We conduct offline experiments on three datasets, with details provided in Appendix 6.
 5.1.2. Baselines
@@
 Text-enhanced models, including P5 (Geng et al., 2022) and PTab (Liu et al., 2022), aim to describe data in natural language to elicit responses from language models.
-CTR-BERT (Muhamed et al., 2021) integrates traditional and textual features for fusion in a MLP, UniSRec (Hou et al., 2022) utilizes a frozen BERT (Devlin et al., 2018) as its text encoder, and CTRL (Li et al., 2023) aligns text and tabular data representations to infuse the ID-based model with semantic information.
+CTR-BERT (Muhamed et al., 2021) integrates traditional and textual features for fusion in an MLP, UniSRec (Hou et al., 2022) utilizes a frozen BERT (Devlin et al., 2019) as its text encoder and CTRL (Li et al., 2023) aligns textual and tabular data representations to infuse the ID-based model with semantics.
+BoW (Zhang et al., 2010) uses bag-of-words representations as item text representations.
 5.1.3. Evaluation Metrics
-In this research, Area Under the Curve (AUC) and logarithmic loss (Logloss) are adopted as evaluation metrics, both widely acknowledged in the assessment of CTR models (Song et al., 2019; Wang et al., 2021; Qu et al., 2016; Guo et al., 2017).
-AUC evaluates the model’s capacity to accurately rank a positive instance above a negative one, thereby indicating its ranking efficacy.
-Logloss, on the other hand, quantifies the degree of correspondence between the predicted and actual CTR, offering a metric for the model’s predictive accuracy.
-5.1.4. Implementation Details
-In our offline experiments, models are implemented using RecStudio (Lian et al., 2023) and Transformers (Wolf et al., 2020).
-We integrate RoBERTa (Liu et al., 2019) to enhance the optimal ID-based model.
-AdamW (Loshchilov and Hutter, 2017) optimizer facilitates all training stages.
-For the DAP stage, $\tau_{1}$ ranges from [0.001 to 0.1] and $\alpha$ from [0.1 to 1.5], with a batch size of 32.
-In the ROMA stage, $\tau_{2}$ is 1, batch size is 128, the language model’s learning rate is 5e-5, and projection units are [128, $f_{i}d$], with $f_{i}$ as item-side feature count and $d$ set at 16.
-The MF2 stage uses a batch size of 1024 and standardized hidden units [128, 64, 64] across ID-based models.
-ID-based models that do not focus on user behavior modeling aggregate user behavior sequences using mean pooling.
+In this research, Area Under the Curve (AUC) and logarithmic loss (Logloss) are used as evaluation metrics for CTR models (Song et al., 2019; Wang et al., 2021; Qu et al., 2016; Guo et al., 2017).
+AUC assesses ranking ability, while Logloss measures predictive accuracy.
 5.2. Overall Performance (RQ1)
-We conduct a comparative performance evaluation of CELA against baselines, with the results detailed in Table 2. We observe that:
-  • Within ID-based models, DIN and DIEN excel on Amazon, because their fine-grained modeling of user behaviors caters to the pronounced reliance on user behaviors.
-Conversely, for Movielens and AppGallery, the dependency on user behaviors appears less critical, resulting in no marked advantage for DIN and DIEN.
-  • CTRL ranks second to CELA in most cases due to its integration of the semantics into the ID-based model, distinguishing it from traditional ID-based models.
-Unlike other text-enhanced models, CTRL retains collaborative signals, enhancing its performance.
-  • Among text-enhanced models, P5 and PTab perform worse than most ID-based models.
-This observation is consistent with results from (Li et al., 2023), corroborating that recommender systems based purely on text generally underperform compared to ID-based models.
-  • CELA demonstrates superior performance across all datasets, thereby validating three core assertions: firstly, the utilization of the PLM extracts intrinsic knowledge within the textual feature; secondly, employing the PLM solely as a textual feature encoder prevents information loss caused by an increase in features and enables the detailed delineation of items; thirdly, the integration of collaborative signals significantly enhances model efficacy.
-This last point is further supported by the superior outcomes observed in the CTRL and UniSRec compared to other baselines.
-Table 2.
-Performance comparison with baselines.
-The best and the second best results are in bold and underlined, respectively.
-denotes the statistical significance of the performance difference with respect to the second-best result at the 0.05 level.
+Table 1. Performance comparison with baselines. The best results are in bold, second-best are in underlined. denotes statistical significance (p ¡ 0.05) compared to the second-best.
 Model Amazon MovieLens AppGallery
@@
 CTRL 0.8954 0.4100 0.8385 0.3655 0.8464 0.4729
+BoW 0.8965 0.4110 0.8371 0.3656 0.8457 0.4729
 CELA 0.8996* 0.3996* 0.8426* 0.3574* 0.8481* 0.4704*
+We conduct a comparative performance evaluation of CELA against baselines, with the results detailed in Table 1. We observe that:
+  • PLMs enhance ID-based models as item textual feature encoders.
+UniSRec and CELA surpass ID-based models on MovieLens and AppGallery.
+This demonstrates that PLMs capture valuable information from item textual features not exploited by collaborative filtering, thereby improving model efficacy.
+  • Purely text-based models typically underperform ID-based models.
+This is evidenced by the performance of P5, and PTab, which are inferior to most ID-based models.
+These findings are consistent with the results reported by (Li et al., 2023), which underscore the limitations of exclusive reliance on semantic modeling.
+  • The integration of semantic and collaborative signals leads to improved performance.
+This is demonstrated by CTRL, UniSrec, and CELA, which excel in most cases.
+The injection of semantics into the ID-based model distinguishes them from traditional ID-based models.
+Furthermore, the retention of collaborative signals makes them outperform other text-enhanced models.
+  • Task-oriented alignment is crucial for PLMs to eliminate the representation discrepancy and better adapt to downstream tasks.
+By aligning the outputs of PLMs and ID-based models, CTRL and CELA outperform UniSRec across three datasets.
+This outcome indicates that a simple whitening process in UniSRec is insufficient to bridge the gap.
+Therefore, alignment is essential to tailor PLMs for recommendation purposes.
+  • Using PLMs solely as item textual feature encoders prevents information loss from feature proliferation.
+Compared to CTRL, CELA performs better in all cases since it extracts textual information from behavior sequences by encoding each item’s textual feature individually, thus avoiding lengthy annotations resulting from including all features of an item.
+In contrast, CTRL transforms tabular data into textual data using a hard prompt template and takes the PLM as a textual data encoder, which fails to provide detailed annotations for items in user behaviors due to token number limitations, leading to information loss.
 5.3. Efficiency Measurement (RQ2)
-To explore the practicality of CELA in industrial scenarios, we measure its training duration for one epoch and inference latency for a single batch on the industrial dataset AppGallery, compared to baselines.
-The training duration is calculated for the most time-consuming stage.
-Results presented in Table 3 reveals that:
-  • Within text-enhanced models, the training expenses of P5, PTab, CTR-BERT, and CTRL are notably high, attributable to their complexity, which is implied by the number of parameters.
-the inference overhead of P5, PTab, and CTR-BERT is also very high for the same reason.
-Although reducing the update frequency can reduce training costs, employing an outdated language model adversely affects recommendation performance.
-Therefore, it is not practical to apply them in real-world scenarios.
-  • UniSRec, by not fine-tuning the language model, maintains training and inference costs comparable to those of DIN.
-However, using a frozen language model limits its efficacy in recommendation tasks.
-This limitation stems from the insufficient alignment of the PLM’s knowledge with the specific requirements of the recommendation domain, resulting in suboptimal integration of textual and non-textual features.
-  • The ROMA stage of CELA necessitates fine-tuning of the language model, but this tuning is conducted at the item level rather than the interaction level.
-Typically, the number of items is significantly lower than that of interactions, leading to minimal training costs, particularly in contexts with a limited number of items.
-The training and inference costs of CELA’s MF2 are on par with those of DIN, making CELA industry-friendly.
 Table 2.
 Comparison of training and inference efficiency between different models on AppGallery.
-Params is the number of trainable parameters, the training time is calculated for the whole process and the inference time is calculated for one batch.
-Model Params (M) Inference (ms) Total (h)
+Params is the number of trainable parameters, the inference time is calculated for one batch and the total training time is for the whole process.
+Model Params (M) Inference (ms) Training (h)
 DIN 0.58 61.38 1.76
@@
 CTRL 102.98 61.38 52.29
-CELA DAP (102.88+0.58) 103.46 / (1.07+1.76) 2.83
-CELA ROMA 102.30 / 0.30
+BoW 1.07 70.48 1.59
+CELA DAP 102.88 / 1.07
+CELA ROMA 102.88 / 2.06
 CELA MF2 0.63 69.64 2.29
-CELA / / 5.42
+CELA / 69.64 5.42
+To explore the practicality of CELA in industrial scenarios, we measure its training duration and inference latency on the industrial dataset AppGallery. Results presented in Table 2 reveals that:
+  • Models utilizing hard prompt templates are impractical for real-world applications due to high training costs or inference overhead.
+This is attributable to their model complexity, indicated by the number of parameters.
+Training models like P5, PTab, CTR-BERT, and CTRL on extensive interaction-level textual data requires significant time, leading to infrequent updates and the use of outdated models, which negatively impacts performance.
+Additionally, the high inference overhead of P5, PTab, and CTR-BERT renders them unsuitable for real-time serving.
+  • UniSRec and CELA maintain training and inference costs comparable to DIN.
+However, UniSRec’s performance is limited due to the lack of fine-tuning of the PLM.
+This limitation arises from insufficient alignment of the PLM’s knowledge with downstream tasks, resulting in suboptimal integration of textual and non-textual features.
+In contrast, CELA fine-tunes the PLM with negligible training expense, achieving superior performance.
+  • CELA is efficient in both training and inference, making it industry-friendly.
+Its low training expense results from fine-tuning the PLM at the item level rather than the interaction level, as the number of items is significantly lower than the number of interactions, with a difference of up to 550 times here.
+Moreover, CELA’s inference cost is comparable to that of DIN as it avoids real-time text processing, utilizing an additional text embedding table that occupies only 87.2MB of storage space.
 5.4. Performance w.r.t. Long-tail Items (RQ3)
@@
 This superior performance is attributed to the incorporation of a PLM that leverages embedded external knowledge for text comprehension.
-Notably, when the popularity falls below 30, the improvements of CELA become particularly pronounced.
-This is because, under such conditions, ID-based models struggle to learn effective embeddings owing to their dependence on historical interactions.
+Notably, when the popularity falls below 20, the improvements of CELA become particularly pronounced.
+This is because, under such conditions, ID-based models struggle to learn effective ID embeddings owing to their dependence on historical interactions.
 5.5. Ablation Study (RQ4)
-To explore the impact of different training processes and components on the overall performance of CELA, we conduct an ablation study with designed several variants:
-  • DIN + RoBERTa: Incorporate RoBERTa as the text encoder and directly undergo the MF2 stage with DIN. It is equivalent to our CELA without the DAP and ROMA stages.
+To explore the impact of different training processes and components on the overall performance of CELA, we conduct an ablation study and the results are presented in Table 3.
+Firstly, we propose a variant to verify that the performance gains of CELA are attributed to its semantics rather than additional parameters:
+  • Scaled DIN: Replacing CELA’s non-trainable text embedding table with a trainable one.
+5.5.1. Domain-Adpative Pre-training
+Domain-Adpative Pre-training (DAP) improves performance by enhancing the PLM’s understanding and expressiveness of dataset-specific texts. We design three variants to verify its effectiveness:
+  • DAP w/o MLM: The MLM loss is removed in DAP.
+  • DAP w/o SimCSE: The SimCSE loss is removed in DAP.
   • CELA w/o DAP: The DAP stage is removed from CELA.
-  • CELA w/o ROMA: The ROMA stage is removed from CELA.
-  • DAP w/o SimCSE: The SimCSE loss is removed in DAP.
-  • DAP w/o MLM: The MLM loss is removed in DAP.
-  • ROMA w/o PLM: A text embedding table encoded by the PLM replaces the PLM in ROMA.
-  • ROMA w/ cos: Utilizing cosine similarity in ROMA.
-  • MF2 w/o text: The text embedding table in MF2 is replaced by DAP’s embedding table from the ID-based model.
-Analysis of Table 4 reveals the following insights:
-  • DIN + RoBERTa yields fewer performance improvements compared to CELA, suggesting that the mere inclusion of textual feature is insufficient for effective information extraction and utilization in recommendation tasks.
-The extent of the performance impact varies with the degree of overlap between the PLM’s corpus and the textual features in the dataset.
-  • Removing the DAP stage from CELA leads to a reduction in performance, as it deprives the PLM of the opportunity to comprehend domain-specific knowledge, and the anisotropy issue compromises the PLM’s output expressiveness.
-Nevertheless, the subsequent ROMA stage aligns the PLM output with the recommendation domain, enabling the recommender system to understand PLM outputs.
-Although the PLM’s knowledge is not perfectly tailored for the domain, it results in performance that still surpasses that of DIN.
-  • The removal of the ROMA stage from CELA also leads to diminished performance.
-While the DAP stage augments the PLM’s capability to comprehend and express the textual feature, a discrepancy between the PLM output space and the collaborative space hinders the recommender system’s understanding of PLM outputs for CTR prediction.
-  • Removing the SimCSE loss from DAP leads to a performance decline, as it overlooks the anisotropy issue present in PLMs, thereby diminishing expressiveness.
-Incorporating SimCSE loss is advisable to enhance the quality of text representations.
-  • Removing the MLM loss from DAP results in diminished performance.
-This is because MLM, serving as a widely adopted pretraining loss, facilitates the swift adaptation of language models to specialized domains.
-DAP w/o MLM performs even worse than CELA w/o DAP on Amazon dataset, suggesting that using SimCSE loss only in pretraining is not ideal due to the inconsistency with the objective of PLMs’ training.
-Originally, SimCSE loss was proposed as an auxiliary objective in pretraining, aiming to regularize pre-trained embeddings’ anisotropic space to be more uniform.
-  • Removing the PLM significantly reduces performance due to less semantic information preservation when using ROMA directly on the text embedding table.
-Directly mapping the text embedding table to the ID-based model’s output space, with only a basic projection head, leads to substantial semantic information loss.
-  • Replacing the dot product with cosine in ROMA results in reduced performance in AppGallery dataset.
-So we leverage dot product as the similarity function, and it’s widely used in recommendation system, such as SASRec (Kang and McAuley, 2018), UniSRec (Hou et al., 2022).
-  • The performance of MF2 w/o text is comparable to that of DIN, but significantly lags behind CELA. This indicates that CELA’s performance improvement benefits from semantic information.
-These observations underscore the importance of both the DAP and ROMA stages in CELA for maximizing the PLM’s potential in recommendation systems, highlighting the critical role of adapting the PLM to domain-specific knowledge and tailoring it for task-specific objectives.
-Table 4. Ablation study. The language model is RoBERTa and backbone is DIN.
+As shown in Table 3, removing either MLM or SimCSE losses, or both, results in performance degradation compared to the full CELA, demonstrating that both losses contribute to performance improvement.
+However, using only the SimCSE loss in pre-training, without the MLM loss, performs worse than no pre-training on the Amazon dataset.
+This indicates that SimCSE loss is more suitable as an auxiliary objective to MLM loss, as originally proposed in (Gao et al., 2021).
+5.5.2. Recommendation-Oriented Modal Alignment
+Recommendation-Oriented Modal Alignment (ROMA) boosts performance by aligning the outputs of the PLM with the ID-based model at the item level.
+We designed three variants to verify its effectiveness and explore what accounts for the enhancement:
+  • CELA w/o ROMA: No alignment is applied, and the text embedding table in MF2 is encoded by a non-aligned PLM.
+  • CELA w/o semantics: The text embedding table in MF2 is replaced by ROMA’s ID embedding table, excluding any semantics.
+  • ROMA w/o PLM: The PLM in ROMA is replaced by a pre-encoded text embedding table, with only the table’s weights being directly updated.
+Removing ROMA results in performance degradation, affirming the indispensability of task-oriented alignment of PLM outputs.
+The performance of CELA w/o semantics is comparable to that of DIN but significantly lags behind full CELA, indicating that CELA’s performance gains stem from the integration of semantics and imply negligible semantic loss during training.
+Moreover, directly applying ROMA to the text embedding table, instead of the original PLM, leads to reduced performance due to less preservation of semantics, evident from the reduced number of parameters.
+This underscores the necessity of PLM in ROMA.
+Table 3. Ablation study. The language model is RoBERTa and the backbone is DIN.
 Variants Amazon AppGallery
@@
 DIN 0.8949 0.4110 0.8437 0.4771
-DIN + RoBERTa 0.8952 0.4127 0.8472 0.4719
+Scaled DIN 0.8951 0.4079 0.8452 0.4747
+DAP w/o MLM 0.8981 0.4047 0.8477 0.4711
+DAP w/o SimCSE 0.8989 0.4033 0.8478 0.4710
 CELA w/o DAP 0.8986 0.4012 0.8475 0.4715
 CELA w/o ROMA 0.8965 0.4050 0.8475 0.4709
-DAP w/o SimCSE 0.8989 0.4033 0.8478 0.4710
-DAP w/o MLM 0.8981 0.4047 0.8477 0.4711
+CELA w/o semantics 0.8959 0.4089 0.8440 0.4765
 ROMA w/o PLM 0.8669 0.5534 0.8463 0.4731
-ROMA w/ cos 0.8997 0.4001 0.8470 0.4715
-MF2 w/o text 0.8959 0.4089 0.8440 0.4765
 CELA 0.8996 0.3996 0.8481 0.4705
 5.6. Further Analysis
-5.6.1. Compatibility for language models.
-We compare the performance of CELA equipped with PLMs of various sizes: BERTTiny (14.5M) (Jiao et al., 2019), BERTSmall (29.1M) (Bhargava et al., 2021), BERTMedium (41.7M) (Turc et al., 2019), RoBERTaBase (110M) (Liu et al., 2019), BERTLarge (336M) (Devlin et al., 2018), and OPT (1.3B) (Zhang et al., 2022).
+5.6.1. Compatibility for language models
+We compare the performance of CELA equipped with PLMs of various sizes: BERTTiny (14.5M) (Jiao et al., 2020), BERTSmall (29.1M) (Bhargava et al., 2021), BERTMedium (41.7M) (Turc et al., 2019), RoBERTaBase (110M) (Liu et al., 2019), BERTLarge (336M) (Devlin et al., 2019), and OPT (1.3B) (Zhang et al., 2022).
 For the OPT model, Causal Language Modeling (CLM) pre-training loss and LoRA fine-tuning (Hu et al., 2022) are implemented.
 Experimental results are illustrated in Figure 4, leading to the following insights:
-  • CELA, when enhanced with a PLM, no matter its size, exhibits significant performance gains over DIN, validating that the use of PLMs can extract knowledge from textual features and benefit CTR prediction.
-  • As the complexity of the comprehensive PLM escalates, the marginal gains in CELA’s performance exhibit a plateauing trend.
-Notably, BERTLarge underperforms relative to RoBERTaBase, attributable to the former’s significantly larger parameter set, which does not fine-tune as effectively with limited items.
-This observation implies that beyond a specific threshold, the addition of parameters ceases to yield proportional improvements, suggesting that a medium-sized PLM is sufficient to attain optimal performance.
-  • CELA with OPT exhibits a decrease in performance compared to comprehensive PLMs.
-This reduction is not only because the extensive parameter space of such models proves challenging to fine-tune effectively with a limited item dataset, but it is also likely due to the generative PLM’s unsuitability as a text encoder for CTR prediction.
+  • CELA, enhanced with a PLM regardless of size, significantly outperforms DIN. This validates that PLMs can effectively extract knowledge from item textual features, benefiting CTR prediction.
+  • As the complexity of the comprehensive PLM increases, CELA’s performance gains plateau.
+Notably, BERTLarge underperforms compared to RoBERTaBase due to its larger parameter set, which does not fine-tune effectively with limited items.
+This suggests that beyond a certain threshold, additional parameters do not yield proportional improvements, indicating that a medium-sized PLM is sufficient for optimal performance.
+  • Generative decoder-based large language models do not enhance CELA as effectively as comprehensive PLMs.
+Specifically, CELA with OPT shows decreased performance compared to comprehensive PLMs.
+This reduction is due to the difficulty in fine-tuning large parameter models with limited data and the generative PLM’s unsuitability as a text encoder for CTR prediction.
 Figure 4. Performance comparison of CELAs with different PLMs on Amazon. The backbone is indicated by N/A.
-5.6.2. Compatibility for ID-based models.
-Given the orthogonal nature of our proposed framework to ID-based models, we explore its efficacy across a variety of backbones, such as AutoInt, DCNv2, DeepFM, and DIN.
-As shown in Table 5, CELA consistently achieves significant performance improvements in comparison to these backbones, thereby confirming the framework’s model-agnostic nature and extensive compatibility.
+5.6.2. Compatibility for ID-based models
+Given the orthogonal nature of our proposed framework to ID-based models, we explore its efficacy and efficiency across a variety of backbones.
+As shown in Table 4 and Appendix C.1, CELA consistently achieves significant performance improvements and comparable inference and training costs when compared to these backbones, thereby confirming the framework’s model-agnostic nature and extensive compatibility.
 Table 4. Performance comparison of models with different backbones. The language model is RoBERTa.
@@
 $\text{CELA}_{\text{DIN}}$ 0.8996 0.3996 0.8481 0.4705
-5.6.3. Alignment of Popular Items Only
-ID-based models exhibit poor performance in cold-start scenarios, suggesting inadequacies in learning item embeddings for cold-start items.
-We aim to ascertain whether aligning poorly learned item embeddings during the ROMA stage might lead to a decline in performance.
+5.6.3. Analysis of Alignment w.r.t. Popular Items
+Figure 5. Performance with alignment w.r.t. popular items, evaluated on Amazon. The $x$-axis represents the top percentile of popular items.
+ID-based models exhibit poor performance in cold-start scenarios, suggesting inadequacies in learning ID embeddings for cold-start items.
+We aim to ascertain whether aligning poorly learned ID embeddings during the ROMA stage might lead to a decline in performance.
 To this end, we design an experiment in which CELA’s ROMA stage aligns only the top 20%, 40%, 60%, 80%, and 100% of items by popularity.
-Figure 5 indicate that CELA’s performance reaches its apex when aligning 80% of the most popular items.
-This suggests that limiting the alignment to a subset of popular items facilitates a more accurate mapping of item text representations to the collaborative filtering space, while aligning poorly learned embeddings will lead to performance degradation.
-Figure 5. Performance with alignment of popular items only, evaluated on Amazon. The $x$-axis represents the top percentile of popular items.
+Figure 5 indicates that CELA’s performance reaches its apex when aligning 80% of the most popular items.
+This suggests that limiting the alignment to a subset of popular items facilitates a more accurate mapping of item text representations to the collaborative filtering space while aligning poorly learned ID embeddings will lead to performance degradation.
 5.7. Online A/B Testing
-To evaluate the effect of CELA on real-world scenarios, we deploy CELA and conduct an online A/B test in an industrial Android AppGallery, which is one of most popular App stores of China.
+To evaluate the effect of CELA on real-world scenarios, we deploy CELA and conduct an online A/B test in an industrial app advertising platform, which is one of the most popular app stores.
 5.7.1. Deployment
-The online deployment process is outlined as:
-  • Data Preprocessing.
-Utilizing a LLM, we refine the collected app descriptions to eliminate extraneous details, such as contact information, thereby enhancing conciseness.
-The refined text is indexed by the item identifier for low-cost storage.
-  • Update Process.
-Given the infrequent changes in app descriptions, weekly training is conducted during the DAP stage.
-Thanks to the efficiency of the ROMA stage, it is feasible to align the PLM daily with the item embeddings from the previous day’s recommendation model.
-This alignment supplies the MF2 stage with the text embedding table, minimizing additional training overhead.
-Accordingly, updates to the MF2 stage also occur daily.
-  • Online Services.
-The online recommendation model retrieves text representations directly from the pre-computed text embedding table via item identifiers, circumventing the need for real-time PLM text processing.
-This deployment facilitates iterative and daily enhancements of both the ROMA and MF2 stages, while the efficient and effective online serving ensures optimal service delivery.
+The industrial app advertising platform, illustrated in Figure 6, comprises two main components: offline training and online serving.
+In offline training, an LLM refines the collected app textual features from the app pool, removing extraneous details such as contact information to enhance conciseness.
+The clean text is indexed by the app identifier for low-cost storage.
+Given the infrequent changes in app textual features, training is conducted weekly during the DAP stage.
+The efficient ROMA stage enables daily alignment of the PLM with ID embeddings from the current ID-based backbone, providing the MF2 stage with the text embedding table and minimizing additional training overhead.
+Consequently, updates to the MF2 stage also occur daily.
+In online serving, when a user initiates a request, the candidate apps are scored by CELA, which retrieves the text embeddings of candidate apps directly from the pre-computed text embedding table, bypassing real-time text processing.
+The candidate apps are then ranked according to Effective Cost Per Mille (eCPM) and the predicted Download-Through Rate (DTR) scores, with the top-$k$ apps displayed in the list.
+Figure 6. Overview of the app advertising platform.
 5.7.2. Online Experimental Results
 Table 5. Performance of CELA in online A/B test.
-DTR Impr eCPM Impr Latency
-0.92% 0.97% 69.73ms
-An online A/B test of the CELA is conducted, utilizing 1% of AppGallery’s extensive traffic—which includes tens of millions of daily active users—over a week.
+eCPM DTR Training Time Inference Time
++1.48% +0.93% -19.41% +9.69ms
+An online A/B test of CELA is conducted on the advertising platform over three weeks, including tens of millions of daily active users.
 The treatment group implements CELA’s three-stage framework, extending the control group.
-Table 6 reveals CELA’s significant impact: a 0.92% increase in Download-Through Rate (DTR) and a 0.97% rise in Effective Cost Per Mille (eCPM), indicating notable improvements in user engagement and revenue.
-Importantly, the online serving latency of CELA, complying with the scenario’s standard of under 100ms and compared to the control group’s 60.04ms, affirms its suitability for real-world applications.
+Each group is trained in a single cluster, where each node contains an 8-core Intel(R) Xeon(R) Gold 6151 CPU (3.00GHz), 32GB RAM, as well as 1 NVIDIA TESLA V100 SXM2 GPU with 32GB memory.
+Table 5 shows the superiorities of CELA: a 1.48% rise in eCPM and a 0.93% increase in DTR, indicating a positive influence on user engagement and revenue.
+Notably, the latency of CELA is below the 100ms threshold , which is acceptable for real-world applications.
+Figure 7. Performance comparison w.r.t. unpopular apps. The bar graph represents the exposure for each group, while the line chart represents the relative improvement compared to the control group.
+To explore the impact of CELA on cold-start items in the real business scenario, we group apps according to baseline exposure and measure the relative improvements in eCPM and DTR.
+Each group consists of an equal number of apps, focusing on the 15 least popular groups as illustrated in Figure 7.
+The results demonstrate that CELA significantly improves performance in unpopular groups.
+This enhancement is credited to the PLM that leverages external knowledge to create more accurate item profiles.
 6. Conclusion
 In this study, we introduced a novel cost-efficient framework for CTR prediction that leverages the strengths of PLMs and textual features alongside ID-based models.
-To tailor the capabilities of PLMs for recommendation, we implemented domain-adaptive pre-training and item-level alignment, concurrently addressing the inefficiencies associated with PLM training at the interaction level.
-By abolishing hard prompts and employing PLMs solely as text encoders, we intricately describe user behaviors while preventing token overflow due to the increase in features.
+To tailor capabilities of PLMs for recommendation, we implemented domain-adaptive pre-training and item-level alignment, concurrently addressing the inefficiencies associated with PLM training at the interaction level.
+By abolishing hard prompts and employing PLMs solely as text encoders, we intricately describe user behaviors while preventing token overflow.
 Our framework is model-agnostic and industry-friendly, with low training overhead and serving latency.
 Comprehensive experiments conducted on both public and industrial datasets validate CELA’s notable enhancements in performance.
+Appendix A Corpus Construction
+For illustration, consider the features of an item detailed as follows:
+    "item_id": 31852,
+    "brand": "Coxlures",
+    "categories": ["Sports & Outdoors", "Dance"],
+    "description": "The ballet skirt is nice for
+                    dress-up play for ballerinas."
+Categorical features such as item_id, brand, and categories can be encoded using conventional methods like one-hot or multi-hot encoding.
+However, textual features like the description, require language models for encoding.
+To enhance the language model’s understanding of dataset-specific texts, a corpus is constructed from item descriptions.
+This corpus is organized as follows:
+    item 1’s description,
+    item 2’s description,
+    item |I|’s description
+Appendix B Experimental Settings
+B.1. Datasets
+Offline experiments are conducted on the datasets below. The statistics are presented in Table 6.
+Table 6. Statistics of datasets for offline experiments.
+Dataset #Users #Items #Features #Samples
+Amazon 300,562 100,682 6 4,494,230
+MovieLens 92,911 26,084 4 10,146,329
+AppGallery 11,056,819 29,759 8 16,398,193
+[…]
+MovieLens Dataset  [^2] https://grouplens.org/datasets/movielens/latest/ collects the rating activity from MovieLens  [^3] http://movielens.org/.
+We extract movie summaries as the textual feature from provided hyperlinks and sample 10 million interactions through random sampling of users.
+The task is to predict whether a user will like the next item, indicated by a rating of 3 or higher, based on a sequence of up to 20 user behaviors.
+AppGallery Dataset comprises application exposure and download records sampled from an industrial app advertising platform over eight days with user consent.
+Data from the first seven days constitute the training set, while data from the eighth day serve as the test set.
+The task is to predict whether the user will download the next app, using app descriptions as the textual feature and a sequence of up to 14 user behaviors.
+B.2. Implementation Details
+In our offline experiments, models are implemented using RecStudio (Lian et al., 2023) and Transformers (Wolf et al., 2020).
+We integrate RoBERTa (Liu et al., 2019) to enhance the optimal ID-based model.
+AdamW (Loshchilov and Hutter, 2017) optimizer facilitates all training stages.
+For the DAP stage, $\tau_{1}$ ranges from [0.001, 0.01, 0.05, 0.1] and $\alpha$ from [0.1, 0.5, 1.0, 1.5], with a batch size of 32.
+In the ROMA stage, $\tau_{2}$ is 1, batch size is 128, the language model’s learning rate is 5e-5, and projection units are [128, $f_{i}d$], with $f_{i}$ as item-side feature count and $d$ set at 16.
+The MF2 stage uses a batch size of 1024 and standardized hidden units [128, 64, 64] across ID-based models.
+ID-based models that do not focus on user behavior modeling aggregate user behavior sequences using mean pooling.
+Appendix C Experimental Results
+C.1. Compatibility for ID-based models
+We extend the evaluation of CELA’s model-agnostic nature by testing it with a wider range of ID-based models and providing a detailed analysis of the computational costs associated with each integration.
+The results are presented in Table 7.
+It is worth noting that CELA’s training involves both the PLM and the backbone, with the PLM requiring 1.07 hours for pre-training and weekly updates, while the backbone is updated daily.
+Table 7. Training and inference efficiency on AppGallery. The inference time is calculated for one batch and the training time is for the whole process.
+Model Inference (ms) Training (h)
+AutoInt 62.83 2.90
+$\text{CELA}_{\text{AutoInt}}$ 77.96 6.99
+DCNv2 59.63 2.28
+$\text{CELA}_{\text{DCNv2}}$ 64.16 5.66
+DeepFM 62.67 2.88
+$\text{CELA}_{\text{DeepFM}}$ 65.36 4.77
+xDeepFM 60.96 2.33
+$\text{CELA}_{\text{xDeepFM}}$ 65.36 4.85
+DIN 61.38 1.76
+$\text{CELA}_{\text{DIN}}$ 69.64 5.42
+DIEN 66.94 2.93
+$\text{CELA}_{\text{DIEN}}$ 75.78 5.86